A Complete Bibliography of The American Statistician: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

11 March 2017 Version 1.01

Title word cross-reference

(X^{2} − Y^{2})^{1/2} [Beh70]. 2^{5} [Hun77]. 2^{k} [SM79]. 2^{m} [Wio72]. 2k − 1 [Tau72]. 2 × 2 [Dro74, GM76, MH71, MH75]. X̄ [Shu73]. F [Gue79, HS76b, Rao76, SM79]. K [Lev79, D’A72, Tau72]. N [Len73, BG76]. P [GP75]. R() [SH76]. R^{2} [Nel76]. S [Joh73, Cra73]. s^{2} [Shu73]. T [Sam75, Eis79, Gue75, Haw75, Rao76, Woo76]. z [Eis79].

-Method [Joh73]. -Minimax [Sam75]. -Notation [SH76]. -Values [GP75].

1 [GM76, Ros79]. 130th [Ano70a]. 131st [Ano71a]. 132nd [Ano72a]. 133rd [Ano73a]. 1969 [Ano70-27]. 1970 [Ano70-28, Ano70-41, Ano71-30, Ano71-38, Bar70]. 1971 [Ano71d, Ano71-42, Ano72u, Cad72, Dem71, Gre71, Mos72]. 1972 [Ano72-27]. 1973 [Ano73-38, Bar73, Lev74]. 1974 [Ban74, Coa74, Gar74, Kem75, McC75, Wel75]. 1976 [God77]. 1977 [Bac77, Duf78, McV78]. 1978 [Mos79b, Nor79, Wic79, You79].

3 [BR79].

4-Deck [MBG75].

Abbreviated [Pos74]. Abstracts [Sea73]. Accuracy [WD77]. Action [Dem75]. Activities [Ano70v, Ano70w, Ano70x, Ano70y, Ano70z, Ano71u, Ano71v, Ano71w, Ano71x, Ano72v, Ano72w, Ano72x, Ano72y, Ano72z, Ano73x, Ano73y, Mar76, MF77]. Actual [GM76]. Ad [Ano76j, Ano77k]. Addenda [Ano70d]. Addition [Bet77]. Adjustment [Joh74b]. Admissibility [DM75]. Advisory [Gol75]. Aesthetic [Doa76]. Affairs [Bro76]. Afraid [NJ73]. after [Ano71-40]. against [Fue72]. Agencies [CEF^{+}70, Gol75]. Agreement [CH79]. Agreement/Categorical [CH79]. Agriculture [PRPM79]. alias [KP79]. Allan [God77]. Allocation [Key79a, Key79b, Nis79, Rob79]. Alternative [Fis73, Gue77c, Han79, Jan73, Les73, Sol75, Har76]. Alternatives [Mal78]. American [CEF^{+}70, Gol75, Mos70, Ano73z]. among [WD77]. Amounts [Add74, Ano74e]. Analogies [Eva73]. Analogues [FW76]. Analysis [Add74, Ano74e, Ans73, Bib73, BF78, BP75, BWSW79, BB71b, CM72, CC74, Fox79, Fri77, KK74, Lee78, MH71, Mee78, NHC79, Ros70, Ros71, Sch71, Sea79, Sha79, SU72, Woo76, Mon79]. Ancillary [Pad77]. Anniversaries [Ano70d, Eis70]. Annotated [Sea79]. Annual [Ano70a, Ano71a, Ano72a, Ano72b, Ano73a, Mer71]. ANOVA [Ano78f, D’A72, HW78, Sir71]. Answer [CJ73]. Answers [RB70, Rub70a, RS70, RAM70, Rub70b, RGH71, Rub71a, RA71, RB71, Rub71b, Rub72a, Rub72b, RS72, Rub72c, Rub72d, RA73, Rub73b, RR73]. APL [Fox79]. Application [GK73a, Gol73, Hou73, Kab70, Mor73, Sim76, Spe74]. Applications [FW76, MHF75, Wal71, Hog79]. Applied [WM72]. Appraisal [RGW70]. Approach [Fis73, LW79, Mak79, RD76, Smi78]. Approaches [Coo72, Sei75, Urq71]. Approximate [GT74, Gue77d]. Approximation [GT71]. Approximations [Gue79]. April [Ano71d]. Arcsine [Mil78]. Arguments [Bar77, Raf70]. Arithmetic [Blo76]. ASA [Ano71-39, Ano74j, BB72, Gol75, Mer71]. Asked [CJ73]. Aspect [Wio72]. Aspects [Coc71]. Assess [BL78]. Assessing [MHF77]. assisted [MH76, NHC79]. Assisting [Mos76]. Associated [Lev79]. Association [Fuc79, Ano73z, CEF^{+}70, Gol75, Mos70]. Assumption [Mun77]. Assumptions [MHL77]. Asymptotic [GK73b]. Attainability [Ano79g, GM77a]. Audio [Pos76]. Audio-Visual [Pos76]. Audiotape [Gri79]. Audiovisual [Pos77b]. Author [Ano79a]. Autocorrelated [Hua76]. Autocorrelation [JM76]. Automated [Lor75]. Autoregressive [Nel76]. Autoregressive-Moving [Nel76]. Available [Gue77e]. Average [Fuc79, Nel76, Sei75]. Averages [Ano78e, Gel78].

B [Ano70-28, Wel75]. Back [Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano74a, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79b, Ano79c, Ano79d, Ano79e]. Bad [Wil78a]. Balanced [Hoc74]. Bands [BPJ73]. Barnes [Coa74]. Baseball [GOS78]. Based [HA75, Tan79]. Basic [Lee78]. Basis [Dem75]. Basu [KT75]. Baye [Ait71]. Bayes [Kru72, TB73]. Bayesian [Dia75, Eas73, EJO76, Lig75, LP76, Nov75, RD76]. Bayesians [Fis72]. Be [Wil78a, BS75]. Beattie [McC75]. Behavior [Bro74a, Fis72]. Behavioral [Kru71]. Being [CJ73]. below [Gue71a]. Benford [Var72]. Bernoulli [Str79, LP76]. Best [Tau72]. Between [D’A72, Guj70a, Luc73, BZ76, Guj70b, JK76, Pro73, Rao76, Wei70]. Bias [Hua76, Ksh71, MHF77]. Biased [DR79, Fuc79]. Bibliography [Pos76, Pos77b]. Bimodality [Hil71]. Binary [Mor78, SU72]. Binomial [Ano73o, Che71, Dro74, Dwa79, GT74, Her72, Rao77, Sch76]. Birnbaum [God77]. Birthday [Rus76, Gli70, Mun77]. Bivariate [Jos70, Kow73]. Bizarre [Bra77]. Blackjack [Ano75e, MBG75]. Blue [Blo76, MHL77]. BMDP [Fra77]. Board [Ano70b, Ano71c, Ano73b]. Book [Car78a, IK78, Noe77, Pos74, Pos79, Shu78b, Shu78a, Sog79, SL79, Tra78]. Both [RW74]. Bounds [Ano79g, GM77a]. Bowman [Mos72]. Box [HMC78, MTL78]. Bridge [Coc71]. brief [BR78a]. Bruyere [Ano71-31]. Buffon [PW75]. Bunches [Fis72]. Bureau [BB71a]. Burgess [You79]. Business [Shi71a].

C [Ano70-27, Bar73, Dem72]. CADA [NHC79]. Calculating [Raf70]. Calculation [Gue77a, Gue77b, Gue79, Hua76]. Calculations [WD77]. Calculators [Gue77e]. Can [Rob78a, Rob78b, Wel78, Wil78a, Win78]. Car [Ten71]. Card [Abr73]. Care [Ric77]. Carlo [LH72, RB74]. Carnival [BR78b]. Case [GM76, Hol74, Sir71, Wio75]. Categorical [SC71]. Cauchy [SB71]. Caused [Fri77]. Census [Bro71, Key79a, Key79b, Nis79, Rob79, Ros71, VK79]. Centers [RM77]. Central [Gue75, Mak79, Mos76]. Certain [BR78b, Jos70, Man72, SB71, WK70, Kap79]. Certified [BS75]. Chain [Con77, Kab70]. Chance [GK79]. Change [Sei75, Har76]. Chapter [Ano70j, Ano70k, Ano70l, Ano70m, Ano70n, Ano71j, Ano71k, Ano71l, Ano71m, Ano71n, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l]. Characteristic [Man73a, Ger79]. Characterize [Nel77]. Charts [Hou73]. Chebyshev [Ano79g, GM77a]. Chi [D’A72, DH72, Gue77d, JB70]. Chi-Square [Gue77d, JB70]. Chi-Squared [D’A72, DH72]. Chicago [Ano71-37]. Chill [Wio75]. Chinese [Orl74]. Choices [Mor78]. Chow [Ano79h, Wil78b]. Civil [Ben72]. Class [Jan71]. Class-Size [Jan71]. Classical [Man73b]. Classification [Spe74, Han79]. Classifications [Doa76]. Classroom [FT79]. Clerk [Bro74b]. Clinical [Ede79]. CLT [Mak79]. Cochran [Fis73]. Coefficient [Bar74, Cro72, Eva73, FT71, Gue77b, Kat75b, Kym70, Win79]. Coefficients [GT74, Guj70b, Guj70a, Han74, LL75, LMR75, Luc73, SU72, Swi74]. Collaborators [Ano72m, Ano77f, Ano78g, Ano79i]. Colleges [MF77]. Combinations [Mor75]. Combinatorics [JK76]. Combined [WM72]. Comment [Coo72, DuM79, Hei79, Nis79, Rob79, Rya79, Wel78, Wel79, Win78]. Comments [Eck72, TB73]. Commission [Eck72, RWM^{+}72, Wei74]. Committee [Ano71-37, Ano76j, Ano77k, Mos70, Rob71, Ros70, Whi71, Ano72-28, Ano73-27, Ano73z, CEF^{+}70, Gif70, Gol75, Mar74, Mar76]. Committees [Gol75]. Common [Bra77, HS76b]. Communication [BB71a]. Community [Ban72b]. Comparing [Adc73, Bro71, MHF75]. Comparison [Bea76, EF77, Har70, RB74]. Comparisons [Lev79]. Compartmental [Kni73, Mat73]. Competitions [GOS78]. Complete [JKS73, Wac76]. Completeness [DM75, Sti72]. Components [Rut71, ST73, SR74]. Compound [Ste73a]. Computation [GM78, HA75]. Computation-Based [HA75]. Computations [RB74, LH72]. Computer [AAF72, BB71a, CS79, DR79, DuM79, Gen77, Hei79, HMC78, MHF75, Mee78, Rya79, Sea79, Sha79, Tan79, Thi79a, Thi79b, WM72, WM73, Wel79, MH76, NHC79]. computer-assisted [MH76, NHC79]. Computer-Based [Tan79]. Computing [Add74, Ano74e, DuM79, Hei79, Hol74, RJ73, Rya79, Thi79a, Thi79b, Wel79]. Concepts [Lee78, Shi71a]. Concerning [Gue77a]. Condensed [CEF^{+}70]. Conditional [Ban71b, Hun72, Law76, Pan73, Str71, Wal71, Kap79]. Conducive [Bra77]. Conduct [MF77]. Conducted [Hun77]. Conference [Ano73z, Ano74j, Shu78b, Shu78a]. Confidence [And72, BPJ73, Des71b, ES79, GT74, Gue71b, Gue72, Gue77a, Hsi72, Kat75a, Noe72, Wil73a]. Confidence-Interval [Hsi72]. Confidentiality [Dun76, Ano77k]. Considerations [FHV75]. Consistency [RW74]. Constructing [Day72, WH71]. Construction [Con77, Jos70]. Consultation [BF78]. Consulting [Ban71a, Boe72, Mar79b]. Consumption [Ken76]. Contestants [Cow75]. Contested [DGK78]. Contexts [Wal76]. Contingency [Fox79, GM76, KK74, Lee78, MH71, MH75, Sne74]. Contrasts [CM72]. Contributions [Fri77]. Control [Nel79]. Convenient [Smi79]. Convictions [Fri72]. Cooperative [Ban71a]. COPSS [Ano72-28]. Correction [Ano74e, Ano76e, Nar74]. Corrections [Ano73n, Ano73m, Ano73o, Ano75e]. Correlation [BS79, Cro72, Eva73, FT71, Gue77b, Kat75b, Kym70, Law76, Law79, LL75, LMR75, Luc73, Mey74, Sch79b, War73, Wei70, Win79]. Correlations [Man75]. Correspondence [Rao76]. Corrigenda [Ano70d, Ano77e, Ano78f, Ano78e, Ano79f, Ano79g, Ano79h]. Corrigendum [And75, Ano73p]. Cost [Wil73a]. Council [Mos70]. Counterviews [Wei74]. Countries [Bro71]. Course [CCF^{+}77, Dia75, Hou73, Lig75, Nel79, Nov75]. Courses [Ban72a, Urq75]. Covariance [Hun72, SR74, Smi78, Wal71, Wil73b]. Coverage [FW76]. Cox [HMC78].

Craps [Bry73]. Criteria [FHV75, Fri72]. Criterion [Ano79f, BA76, Eat78, Pol78]. Critical [Joi74]. Crooked [Bry73]. Crowder [Bar70]. Cuba [Gli70]. Cumulants [Ano73n, Dos73, Har72b]. CUPM [Ano72-39]. Curio [GD74]. Curious [Wio72]. Current [RM77]. Curriculum [Pie76]. Curve [BPJ73]. Cycle [Shi71a].

Data [Bro71, CH79, Dav76, Day72, Fre71, Joi74, Men70, Mor73, Nam78, NHC79, Ros71, Sea79, SH76, VK79, Wil78a, WH71]. Dated [Gli70]. Davis [Wil76]. Decision [Har72a, Woo76]. Deck [MBG75]. Decomposition [Pan73]. Decompositions [DH72]. Deficiencies [Kot75]. Degree [Sei75, Sir71]. Degrees [Goo73]. Demographic [RM77]. Denominator [HS76b]. density [TK76]. Department [Ano71-37]. Dependence [Han71, HS76b]. Dependent [YY75]. Derivation [Joh73, Mar70, Whi73, Win79]. Deriving [Wac76]. Design [FF73, Had73, Hun77, MHF77, PRPM79, RGW70, Scl72]. Designed [Har74]. Designs [Coc71, MHF75]. Desk [Gue77b, Gue77e]. Determination [Bar74, Gue70, Gue77c]. Develop [BL78]. Developing [Dun77, Ede79]. Development [CCF^{+}77, Kat75b, Ric77]. Developments [Shi71b]. Deviance [Arn74]. Deviation [D’A70, GT71]. Device [Pos78b]. Diagrammatical [Hsi72]. Diagrams [Ryb75]. Dice [Bry73, DD72]. Dichotomous [D’A72]. Didaktik [Noe77]. Difference [And72, BZ76, JB70]. Differences [Luc73]. Differentials [Gas75]. Digest [Bry76]. Dimension [AM74, And75]. dimensional [Mon79]. Direct [Kat75b]. Directors [Ano70b, Ano71c, Ano73b]. Discrete [Sel71, Whi73, KP79]. Discrimination [Fue72, KA75]. Discussion [Dia75, Hoc73, Lig75]. Display [Sne74]. Displays [Wai74]. Distributed [Coo72]. Distribution [Abr73, Ano73o, Bra77, Che71, Con77, D’A70, Dwa79, GK73b, Gue70, Gue71a, Gue72, Gue77b, Gue78b, Haw74, Her72, Hoc74, JB70, KS72, Noe72, Rao77, Tau72, Ger79, Kap79, KP79, Tra78]. Distribution-Free [Gue70, Noe72]. distributional [Kap79]. Distributions [Ban71b, EJO76, Jan71, Jos70, KJ73, Kow73, LB71, Man72, Man76a, Nel77, Pro73, RW74, Whi73]. Disturbances [Hua76]. Do [DD72]. Don [HR71]. Donald [Bar73]. Double [DF77, Mil78]. Dream [Orl74]. Dual [DF77]. Due [Ksh71]. Dummy [GD74, Guj70b, Guj70a, Men71].

Each [Max74]. Early [Mos79a]. Earnings [Gas75]. Easily [Gue78a]. Easy [Wat73]. EASYMAP [IS79]. Economic [Ano70p, Ano70o, Ano70-39, Ano70-40, Ano71b, Ano71o, Ano71-40, Ano71-41, Ano72b, Ano72o, Ano72n, Ano72-40, Ano72-41, Ano73r, Ano73q, Ano73-39, Ano73-40, Bar77, BB71a, Joh74b, Mal78]. Editor [Ano73m, BMF^{+}72, BDR^{+}72, CTL^{+}77, CN78, CNR^{+}72, DSMG78, DTB^{+}70, DMK^{+}70, DWTG78, DGK^{+}75, DDB^{+}75, FGF^{+}71, FvFC^{+}74, FDH^{+}73, GRB^{+}74, GDS^{+}76, GFM^{+}77, HPG^{+}71, HJN^{+}73, HHN^{+}76, HGS^{+}70, KBE^{+}73, KBR^{+}71, KBT^{+}75, KsLG^{+}74, LFB^{+}72, MDH^{+}71, MRT^{+}79, MDC^{+}72, Nar74, NVM^{+}78, PKT^{+}79, PCM^{+}74, PGG^{+}79, RWH^{+}70, SGM77, SHG76, SBK^{+}75, SZM79, SKR^{+}71, Var72, WSG^{+}73, WRH^{+}70, WER^{+}77, YYB^{+}76]. Editorial [Ano72m, Ano77f, Ano78g, Ano79i, WC70]. Education [Car78b, Cle71, CEF^{+}70, Fau77, Fed78, Hog72, Pos76, Pos77b]. Edward [Bar70]. Effect [MHL77, Rus76]. Effective [Joh74a]. Effects [CM72]. Efficiency [MHL77]. Efficient [Han74]. Eisenhart [Kut74]. Election [DGK78]. Elections [RS73, She72]. Elementary [Gil77, Hoy72, Mar79a, Whi73]. Empirical [Kru72, Wai74]. Employees [Man74]. Employment [Fue72]. Energy [Mos79a]. Eng [Car78a]. Engineering [SS70]. Enough [GOS78]. Environmental [Fri77]. Equality [D’A72, EF77, Guj70b, Guj70a, SC71]. Equally [DD72]. Equations [SC71]. Equivalence [Sol75]. Equivalency [KA75]. Ernest [Mos79b]. Error [Bli71, Blo78, BZ76, Des71a, FT79, GM76, Bar76]. Errors [AM74, And75, Fei71, Fri72, GK73a, Lor75, Rao73]. Errors-in-Variables [Rao73]. Establishing [Ban71a]. Estimability [See77]. Estimate [Gue71a, Kat75a]. Estimates [Arn72, Dea70, GT74, LB72, Woo76]. Estimating [Max74, Men71, Rub73a, SR74, Han79]. Estimation [Bha76, Bli71, Bus73, Che71, Coo72, D’A70, Eas73, GT71, Hsi72, Kru72, Sam75, Scl72, Sel71, Smi78, Sti72, Str79, YY75, TK76, Tor78]. Estimators [Bha73, Gue78a, Han74, Hua76, MHL77, ST73, Sch76, TB73]. Ethical [Bra73]. Ethics [Gib73]. Evaluating [CCF^{+}77, Cow75, MHF75]. Evaluation [DN73, FHV75, Joi74, MHF77, Ano71-37]. Evenings [Coc71]. Events [Won72]. Evidence [Kat75a]. Exact [MBG75, Man73b, SM79]. Example [Ano79f, Bro74a, Dea70, FT79, Har74, JM76, Lee78, Pol78, Rao77]. Examples [Ano78e, Arn72, Cut76, DH72, Gel78, Hun77, Mal79]. Excerpts [Kru71]. Exchangeable [HS76a]. Exercise [BP75, Man73a]. Exercises [Nam78]. Exist [DD72]. Expansion [RLM74]. Expect [Ban70]. Expectation [Beh70, Hun72]. Expectations [Bet77, Gol73]. Experiment [Har74, SM79]. Experimental [AM74, And75, Had73, Har75, MHF75, MHF77, PRPM79, Scl72]. Experiments [Add74, Ano74e, Hun77]. Expert [VC76]. Explaining [Bro76]. EXPLOR [MHF77]. Exploratory [LW79]. Exponent [Ano73o, Her72]. Exponential [Pro73, Sel71]. Expository [Coo72]. Expressing [Ano73n, Har72b]. Expression [Jan73, Les73]. Extended [Sea73]. Extensions [WS79]. Extreme [Arn74]. Ezekiel [Wel75].

Factor [Sch71, Mon79]. Factors [Add74, Ano74e, Gue77a]. Failures [Ste73a]. False [Har74]. Family [Sel71]. Fatou [Sim76]. Features [MHF77]. Federal [Ano73z, Dun77, Eck72, Gol75, Mal78, RWM^{+}72, Shi70, Shi71b, Wei74, Ano73z, Ano76j]. Feller [Ano70-41]. Fieller [Zer78]. Films [KK78]. Filmstrips [Gri79]. Finding [Key70]. Findings [BB72]. Finetti [HS76a]. Finite [Bea76, BZ76, Kab70]. First [Ano70o, Ano71o, Ano72n, Ano73q, Len73]. Fisher [Bea76, Win79]. Fit [Bus73]. Fixed [CM72, Des71b]. Food [Joh74b]. Forced [Mor78]. Forecast [Des71a]. Forecasting [MHW74, MW78]. Form [Boo75, HMC78, Joh73]. Forms [Beh77]. Formula [Ano73p, Gok72, Gue75, Man75, Wal71]. Found [Gue78a]. Foundations [God70, HR71, Sav70]. Four [Ryb75]. Fourth [Ano70p, Ano71b, Ano72o, Ano73r, Shu78a]. Fraction [Gue71a]. Framework [CCF^{+}77, Dun77]. Franklin [Dem71]. Frederick [Dem71]. Free [Gue70, Noe72]. Freedom [Goo73, Sir71]. Freeman [Mil78]. Frequency [Doa76, EJO76]. Front [Ano70q, Ano70r, Ano70s, Ano70t, Ano70u, Ano71p, Ano71q, Ano71r, Ano71s, Ano71t, Ano72p, Ano72q, Ano72r, Ano72s, Ano72t, Ano73s, Ano73t, Ano73u, Ano73v, Ano73w, Ano74f, Ano74g, Ano74h, Ano74i, Ano75f, Ano75g, Ano75h, Ano75i, Ano76f, Ano76g, Ano76h, Ano76i, Ano77g, Ano77h, Ano77i, Ano77j, Ano78h, Ano78i, Ano78j, Ano78k, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n]. Function [Gil77, Jan73, Les73, Man73a, Raf70, RLM74]. Functional [Beh77, HMC78]. Functions [Hoy72, Kot75]. Future [Ano71-39, Dun76, Fol70].

Gainsbrugh [Bac77]. Game [BR78b, CC74, GK79, Yeh70]. Games [GOS78, PRPM79]. Gamma [Gue72, Raf70]. General [Bli71, Hoc74, Nor76, Zer78]. Generalization [Guj70a, Har76]. Generalizations [Gin71, Kea74]. Generalized [Bha76, Bus73, CM72, Dwa79, Rid77, Sol75, Spe74]. Generating [Hoy72, Jan73, Kot75, Les73, MSH76, RLM74, KP79]. generation [LH72]. Geometric [LL75, Pro73, Sch79b]. Geometrical [Ait71]. George [Ano72u, Ban74, Kem75]. Get [CJ73]. Given [Con77, Gue77c]. Glance [IK78]. GLIM [BR79]. Goals [Ano71-39]. Good [Wil78a]. Goodness [Bus73]. Gosplan [Gif70]. Government [Ano76j, Man74]. Governmental [Sch79a]. Graduate [Ano72-39]. Graphical [Fie79, HM74, Mul72, Sne74, Wei70]. Graphics [Gen77, BR78a]. Graphs [Ans73]. Grouped [Fre71].

H [McC75]. Hamlet [Spr70]. Harold [Lev74]. Health [CEF^{+}70, Fri77, Ric77, Ste73b]. Henry [Cad72, McV78]. Herbert [Duf78]. Heteroskedasticity [JM76]. High [Pie76]. Higher [Fau77]. Highlights [Fra77]. Hijacking [Gli70]. Histograms [DF77, Pos77a]. history [BR78a]. Hoc [Ano76j, Ano77k]. Hockey [Mul77]. Homogeneity [GM76]. Horse [RD76]. Horse-Racing [RD76]. Hotelling [Lev74]. Human [Ano74j, Bro76]. Hypergeometric [Bea76, CS79, Dro74, GT74, Gue78b, Jan73, Les73]. HYPERSAMS [CS79]. Hypotheses [Gue77c, See77, Tau72]. Hypothesis [Hsi72, Kut74, Lor75, Sol75].

IASPS [MN71]. Ideas [Hun77]. II [Fei71, Fri72]. Illustrate [Day72]. Illustrated [Swi74]. Illustrating [Mak79]. Illustration [Mul72]. Immigration [Mat73]. implementation [TK76]. Impossible [Orl74]. Improper [Had73]. Improvement [Mal78]. Improving [Ano76j, Dav76].

Income [Man76b]. Incomplete [Gue72]. Independence [Hun72, Seb73, Shu73, Str71, Wil73b]. Independent [Ano78e, Blo78, D’A72, Gel78, IS79, KT75, Lev79, Ste73a, Won72, Bar76]. Index [Ano79a, Wio75]. Indexes [Men71]. Indistinct [Mor75]. Industrial [Gif70]. Industry [Mar79b]. Inequality [GM77b, WS79]. Inference [Ano79f, Eas73, HM74, LP76, Pol78]. Inferences [Fuc79]. Information [Ano77l, Ano78l, Ano79o, Key79a, Key79b, Mos79a, Nis79, Rob79, RGW70, Sal73]. Innovations [BB71a]. Insanity [Yeh70]. Instability [Swi74]. Instant [Yeh70]. Instruction [WM72, WM73]. Instructional [Tan79, Urq71]. Instrumental [Gin71]. Instruments [CCF^{+}77]. Integers [Len73]. Integral [FW76, Haw75, Hig75, Raf70]. Integrity [Ano73z]. Interaction [Boe72]. Interactions [Har70, Mon79]. Interactive [Gen77, Gre77, MHW74, MW78, Mak79, Sha79, WM73]. Interest [Ano70d, Eis70]. Interesting [Sim76]. International [Ano70v, Ano70w, Ano70x, Ano70y, Ano70z, Ano71u, Ano71v, Ano71w, Ano71x, Ano72v, Ano72w, Ano72x, Ano72y, Ano72z, Ano73x, Ano73y]. Interpretation [BB71b, Chi70, Eva73, EJO76, GP75, Nel76, SU72]. Interpretations [Cro72]. interpreting [Mon79]. Interval [GT74, Haw74, Hsi72, Kat75a, Woo76]. Intervals [Des71b, ES79, Gue71b, Gue72, Gue77a, Noe72]. Introduction [GOS79, TK76, Gri79]. Introductory [CCF^{+}77, Urq71]. Invariance [Bha73]. Invariant [Pad77]. Inverse [BPJ73, CM72, Mil78, Spe74]. Involving [Add74, Ano74e, Tau72, YY75]. Irene [Coa74]. Isador [Wic79]. ISI [Shu78b, Shu78a]. Issue [Car78b]. Issues [Fau77, Ste73b].

J [Ano71-42, Wel75]. J. [Gre71]. Jacobian [Wat73]. Jerome [Ano70-27]. John [Ano70-28, Ano72-27]. Johnson [Ano71-38]. Joint [Ano70a, Ano71a, Ano72a, Ano73a, Ano76j, Mos70]. Jointly [Men71]. Joseph [Wil76]. Jr [McV78]. Judicial [Fei71]. Julius [Nor79]. Jury [Fri72].

Kaysen [Ros70]. Keller [She78]. Kit [Ros79]. Knockout [Wio72]. Known [Bry72]. Kolmogorov [DN73, GK73b]. Kurtosis [Chi70, Dar70, Hil71].

L [Ano71-30]. Lab [Pos78a]. Lack [Ano77e, Cra73, SS76]. Lagrange [NJ73]. Lags [Coo72]. Lansing [Ano70-28]. Large [Ano78e, BG76, BPJ73, Bro74a, Gel78, Luc73]. Laurence [McV78]. Law [Bet77, Fue72, Var72]. Laws [Ano78e, Gel78]. League [Mul77]. Leap [Rus76]. Least [And72, Bha76, Bus73, KBG75, Rid77, SS70]. Least-Squares [SS70]. Lecturer [Sid74]. Legal [Kat75a]. Lehmann [Joh74a]. Lemma [Bry72, KT75, Sim76]. Letter [Var72]. Letters [Ano73m, BMF^{+}72, BDR^{+}72, CTL^{+}77, CN78, CNR^{+}72, DSMG78, DTB^{+}70, DMK^{+}70, DWTG78, DGK^{+}75, DDB^{+}75, FGF^{+}71, FvFC^{+}74, FDH^{+}73, GRB^{+}74, GDS^{+}76, GFM^{+}77, HPG^{+}71, HJN^{+}73, HHN^{+}76, HGS^{+}70, KBE^{+}73, KBR^{+}71, KBT^{+}75, KsLG^{+}74, LFB^{+}72, MDH^{+}71, MRT^{+}79, MDC^{+}72, Nar74, NVM^{+}78, PKT^{+}79, PCM^{+}74, PGG^{+}79, RWH^{+}70, SGM77, SHG76, SBK^{+}75, SZM79, SKR^{+}71, WSG^{+}73, WRH^{+}70, WER^{+}77, YYB^{+}76]. Levels [And72]. Li [Ano70-27]. Life [BWSW79, Key70]. Light [RW74]. Likelihood [Dea70, Oli72, ST73, Sel71, Smi78, Sol75, Han79]. Likely [DD72]. Limit [Bar77, Gue71a, Mak79]. Limitations [Bar74]. Limits [Gue70, Sas70]. Line [IS79, LB71]. Linear [Boo75, D’A70, Fox79, Guj70b, Guj70a, Kut74, Man73b, ST73, See77, Zer78]. Linearity [Mey74]. Linking [Dro74]. List [Eis75, Pos74]. Literary [Bry76]. Literature [KJ73]. Living [Pos77a]. Local [Blu75]. Location [Gre77]. Log [Fox79]. Log-Linear [Fox79]. Logarithmic [YY75]. Logistic [BPJ73]. Loglinear [KK74]. Lognormal [Nel77]. Look [ES79, KBG75]. Low [Haw74]. Lubin [Wic79]. Lucas [McV78].

Machine [IS79, LH72]. Machine-generation [LH72]. Machine-Independent [IS79]. Made [Wat73]. Mahalanobis [Dem72]. Maintaining [Ano73z]. Majorities [RS73, She72]. Make [Key70]. Makers [Mos76, Woo76]. Making [Bry76]. Management [Kab70]. Manual [Car78a]. Manufacturing [Nel79]. Mapping [IS79]. Marginal [Ban71b]. Marginals [Kow73]. Markov [Con77, Kab70]. Marriageable [Rub73a]. Marshall [Duf78]. Martin [Bac77]. Masters [Car78a]. Matching [Abr73]. Materials [Pos76, Pos77b, VK79]. Mathematic [Noe77]. Mathematical [Kru71]. Mathematics [Mos70]. Matrix [Ano78f, HW78, Smi78]. Matter [Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70q, Ano70r, Ano70s, Ano70t, Ano70u, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71p, Ano71q, Ano71r, Ano71s, Ano71t, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72p, Ano72q, Ano72r, Ano72s, Ano72t, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano73s, Ano73t, Ano73u, Ano73v, Ano73w, Ano74a, Ano74b, Ano74c, Ano74d, Ano74f, Ano74g, Ano74h, Ano74i, Ano75a, Ano75b, Ano75c, Ano75d, Ano75f, Ano75g, Ano75h, Ano75i, Ano76a, Ano76b, Ano76c, Ano76d, Ano76f, Ano76g, Ano76h, Ano76i, Ano77a, Ano77b, Ano77c, Ano77d, Ano77g, Ano77h, Ano77i, Ano77j, Ano78a, Ano78b, Ano78c, Ano78d, Ano78h, Ano78i, Ano78j, Ano78k, Ano79b, Ano79c, Ano79d, Ano79e, Ano79j, Ano79k, Ano79l, Ano79m]. Matter [Ano79n, Rob78a, Rob78b, Wel78, Win78]. Maximizing [Joh73]. Maximum [Dea70, Oli72, ST73, Sel71, Smi78]. May [Ban70]. Mean [Bli71, Blo76, BZ76, Cut76, GM77b, HS76b, Man75, Mar70, Sir71, Woo76]. Meaning [Sir71]. Means [Adc73, BZ76, Nor76, PH77, SM79]. measured [Bar76]. Measurement [Blo78, CCF^{+}77, FT79]. Measurements [Max74]. Measures [Gas75, Hil71]. Measuring [Sei75]. Mechanical [Eva73]. Median [GM77b, Lev79]. Medical [She78]. Meeting [Ano70a, Ano71a, Ano72a, Ano73a]. Meetings [Ano71d, Ano72-28, Mer71, Ano70a, Ano71a, Ano72a, Ano73a, Ano73-38]. Membership [Mer71]. Memory [MBG75]. Method [Cow75, Day72, Had73, Joh73, Joh74a, Men71, Shu73, SM79, KP79]. Methodology [GP75, MH76]. Methods [BL78, Bly72, Fie79, Mal79, Nel72, RB74, Han79]. Meyer [Ano71-30]. Midrange [Bro74a, RW74]. Minimal [Ano77e, Joh74a, SS76]. Minimax [Sam75]. Minimization [WK70]. Minimum [Arn72, Blu75, Eat78, Gue71a, Gue78a]. Minitab [RJ73]. Minutes [Ano72-28, Ano73-27]. Missing [Ksh71, Scl72, Sir71]. Misspecification [Rao71]. Mixed [CM76, Hoc73]. Mode [GM77b, Nel77]. Model [CM76, Gas77, Hoc73, Kni73, Kut74, Man73b, Mor78, Rao73, Rut71, ST73, Sch79b, Sch71, Zer78, Mon79]. Models [Cra72, GT74, Hoc74, JK76, KK74, Kut74, Mat73, Nel76, Scl72]. Modernizing [Shi71a]. Modified [Sch76]. Moment [Jan73, Kot75, Les73, RLM74]. Moments [Abr73, Ano73n, Dos73, Har72b, LB71, Pan73, SB71, WD77]. Monitor [NHC79]. Monotonicity [Gil77]. Monte [LH72, RB74]. Mordecai [Wel75]. Motivation [BB75]. Moving [Nel76]. Multi [Ban71b]. Multi-Variate [Ban71b]. Multicollinearity [GK73a, Sas70]. Multidimensional [Lee78]. Multinomial [Nel72, Tor78]. Multiple [Cra72, Cro72, Des71a, Har70, Law79, MSH76, Rao71, Rao76, Wei70]. Multipliers [NJ73]. Multivariate [Bib73, Han79]. Murph [Car78a]. Mutually [Ano78e, Gel78, Won72]. Myth [Bry76].

Name [JKS73]. National [Mar74, Mar76, BB71a, Mos79a, Mos70, Mul77]. Natural [Rao77]. Needle [PW75]. Needs [Har75]. Negative [Ano73o, Her72, ST73]. News [Ano70-29, Ano70-30, Ano70-31, Ano70-32, Ano70-33, Ano70-34, Ano70-35, Ano70-36, Ano70-37, Ano70-38, Ano71y, Ano71z, Ano71-27, Ano71-28, Ano71-29, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano72-29, Ano72-30, Ano72-31, Ano72-32, Ano72-33, Ano72-34, Ano72-35, Ano72-36, Ano72-37, Ano72-38, Ano73-28, Ano73-29, Ano73-30, Ano73-31, Ano73-32, Ano73-33, Ano73-34, Ano73-35, Ano73-36, Ano73-37]. Neyman [Sol75]. Nicholson [Ano72u]. Ninetieth [Nel77]. Nixon [Ano71-40]. No [Ros79, LB71]. Non [Ano79g, GM77a, Gue75, HM74, Kow73, Men71, Raf70, ST73, Sol75, SS70]. Non-Attainability [Ano79g, GM77a]. Non-Central [Gue75]. Non-Equivalence [Sol75]. Non-Integral [Raf70]. Non-Negative [ST73]. Non-Normal [Kow73]. Non-Parametric [HM74]. Non-Seasonal [Men71]. Non-Typical [SS70]. Noncentral [Gue79, Haw75]. Nonlinear [Ano76e, Gal75]. nonparametric [TK76]. Nonstandard [Lor75]. Nonstatistical [Wal76]. Nonverbal [Urq71]. Normal [Adc73, Ban71b, D’A70, Gue71a, Gue72, Gue77a, Haw75, Hol74, Kow73, Seb73, Tau72, Han79]. Normality [MHL77]. Notation [SH76]. Note [Beh77, Coo72, Cra73, FT71, GK73b, Gro71, Gue71a, Guj70b, Har72a, Haw74, Jan71, KA75, LB72, LL75, Mun77, Pro73, Rut71, SR74, SC71, Sol75, Str71, Win79, Won72, Tor78]. Notes [Ano70j, Ano70k, Ano70l, Ano70m, Ano70n, Ano71j, Ano71k, Ano71l, Ano71m, Ano71n, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano73h, Ano73i, Ano73j, Ano73k, Ano73l, Rao71, Rao73]. Null [Sol75]. Number [Mar70]. Numbers [Ano78e, Gel78, Sea79]. Numerian [SL79]. Numerical [Joi74].

O [MHL77]. O-BLUE [MHL77]. Objective [Bly72]. Objects [Mor75]. Observations [Cra73, Fre71, Han79]. Observed [Key70]. Obtain [Gue72, Kap79]. Obtaining [Joh74a, RB74]. Oddity [Oli72]. Odds [MH75, Spr70]. Office [Mos76]. Official [Har72a]. OLS [Hua76]. One [GK73b, Max74, Spe74, Bar76, Hog79]. One-Sample [GK73b]. One-Way [Spe74]. Operational [Ben72, Sch71]. Operator [Goo79]. Optimal [Des71b]. Optimum [MBG75]. Order [Beh77, RB74, SB71, LH72]. Ordered [Gol73, Mar70]. Organization [MF77]. Oriented [Mee78]. Orthogonal [CM72]. Other [Eva73, Wai74]. Outer [PH77]. Outlook [Ano70p, Ano70o, Ano70-39, Ano70-40, Ano71b, Ano71o, Ano71-40, Ano71-41, Ano72b, Ano72o, Ano72n, Ano72-40, Ano72-41, Ano73r, Ano73q, Ano73-39, Ano73-40]. Output [Sea79]. Overall [And72].

P. [Dem72]. Package [BWSW79, Tan79, MW78]. Packages [DuM79, FHV75, Hei79, Rya79, Thi79a, Thi79b, Wel79]. Pairwise [Lev79]. Panel [Ano72-39, Kru71]. Paradox [Kni73, Mat73]. Parameter [Che71, Sch76, Sel71]. Parameters [Man73b]. Parametric [HM74]. Partial [FT71, Fuc79, Kym70, Law76, Law79, LMR75, Luc73, Mul72, War73]. Pascal [Ano73o, Her72, Kea74]. Path [LMR75]. Paul [Ano71-31, Ano71-30]. Peakedness [Dar70]. Pearson [Sol75]. Pedagogical [JM76]. Pedagogy [WM72]. Categorical [CH79]. RUNNER [MW78]. Pennsylvania [Ano71d]. Percentiles [Nel77]. Perpendicular [Mar79a]. Personal [Ano70-34, Ano70-35, Ano70-36, Ano70-37, Ano70-38, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano72-34, Ano72-35, Ano72-36, Ano72-37, Ano72-38, Ano73-33, Ano73-34, Ano73-35, Ano73-36, Ano73-37, Boe72]. Personalizing [She78]. Perspective [Bir71]. Piecewise [Boo75]. Pilot [BL78]. Pioneer [Ban74]. Plan [She78]. Planes [Gli70]. Planning [Dun77, Gif70]. Played [GOS78]. Plot [Gen77]. Plots [Ksh71, MTL78, Mon79]. Pluralities [RS73, She72]. Point [Che71, Eas73]. Poisson [Dro74, Ger79, Mat73, Mul77, Ste73a]. Policies [Ano71-40]. Policy [Dav76, Mos76, Ric77, WC70]. Policy-Relevant [Dav76]. Politics [Hau73]. Poll [Bry76]. Pollutants [Fri77]. Polya [JK76, KS72]. Polynomial [BS79, Coo72, Sei75]. Pooling [LB72]. Poor [Dav76]. Population [Bea76, Joh74b, Kym70, Rub73a, Str79, SL79]. Populations [BZ76, D’A72, Tor78, Ano74j]. Positive [LB71]. Possible [Fuc79]. Posterior [Hol74]. Power [Gil77, Gue77d, Gue77c, Gue79, RLM74]. Powers [Len73]. Practical [Gre77]. Practice [MS75]. Practices [BL78]. Prediction [Rid77]. Predictive [EJO76]. Preliminary [Ano73-38]. Preparation [Ano72-39]. Preposterior [Woo76]. Present [ES79]. Presentation [Kat75a, Nam78]. President [Eck72, Ano71-40, RWM^{+}72, Wei74]. Presidents [Ano72-28, Ano73-27]. Principles [Men70, RGW70]. Printer [IS79]. Prior [Fol70]. Privacy [Ano77k, Ban72b]. Probabilistic [Gil77]. Probabilities [Ano75e, BR78b, Fol70, GM76, Gue77b, Gue77e, Key70, MBG75, Str71]. Probability [Dem75, DGK78, FW76, Gas77, Gri79, Hoy72, JK76, Mor78, Pos78b, Pos78a, Ryb75, Tan79, Yeh70]. Problem [BB75, ES79, Jan71, RD76, Rus76, SS70, Ten71, WK70, Whi71, Gli70, Mun77]. Problems [Fri77, Goo72, Lor75, Mal78, Ryb75, Tem73]. Procedure [And72, KBG75, Str79]. Procedures [Har70, KA75, Wil73a]. Proceeding [Kat75a]. Proceedings [Men79, Shu78b, Shu78a]. Process [Fei71, Har72a, Ste73a, LP76]. Processing [Men70]. Proctorial [WM72, WM73]. Production [Ken76]. Professional [Ano73z]. Program [Ano73-38, CS79, DR79, FHV75, Gre77, HMC78, MHF75, MHF77, Mee78, Sha79, Sid74]. Programmed [MH71]. Programs [KBG75, WD77]. Progress [Mos70]. Projections [Mar79a]. Proof [Blu75, Bry72, Des71a, Gil77, Har76]. Proofs [Bib73]. Properties [Ano73o, Her72, MHF77]. Property [FW76, Man72, Ger79]. Proportions [EF77, Nel72]. Proposal [WC70]. Prospects [KJ73]. Protection [Ban72b]. Protocol [Ede79]. proxy [Bar76]. Public [Bro76, CEF^{+}70, Mos76, Ste73b]. Publications [BB72, Eis75, FF73]. Publishing [Sea73]. Pyramid [Gas77].

Quadratic [Joh73]. Qualitative [Add74, Ano74e]. Quality [Nel79]. Quantiles [Gue77c]. Quantitative [Add74, Ano74e, BR78a]. Quarter [Ano70p, Ano70o, Ano70-39, Ano70-40, Ano71b, Ano71o, Ano71-41, Ano72o, Ano72n, Ano72-40, Ano72-41, Ano73r, Ano73q, Ano73-39, Ano73-40]. Question [CJ73, Gib73]. Questions [Bra73, Car78a, RB70, Rub70a, RS70, RAM70, Rub70b, RGH71, Rub71a, RA71, RB71, Rub71b, Rub72a, Rub72b, RS72, Rub72c, Rub72d, RA73, Rub73b, RR73]. Quixote [HR71]. Quizzing [WM73].

R [Ano70-27]. Racing [Bai77, RD76, Ten71]. Randolph [You79]. Random [Ano78e, Beh70, Gel78, Han71, JB70, MSH76, RB74, KP79]. Randomization [Gre77, Had73, Har75]. Range [Hoc74]. Rank [Eva73, Sch79b]. Ranking [GOS79]. Rapprochement [JK76]. RATCAT [CH79]. Rater [CH79]. Rates [Key70, Sei75]. Ratio [Gue77a, Nel72, Sol75, Han79]. Ratios [HS76b, MH75]. Raymond [Mos72]. Reading [Sav70]. Real [LB71]. Really [Dar70, JKS73, Wil78a]. Recommendations [Har74, Wei74]. Recourse [Ano73o, Her72]. Redeeming [Coc71]. Reduced [Bli71]. Regional [Ano71d, Ano73-38]. Regressing [Han74]. Regression [Ano76e, Ano78f, Arn74, BA76, Boo75, BS79, Cra72, Cut76, Des71a, DR79, Gal75, Han74, HW78, Kym70, LB72, LW79, Man73b, MS75, Men71, Mul72, SC71, SU72, Swi74, Wei70]. Regressions [Fre71, Guj70b, Guj70a, Rao71, Rao76, YY75]. Regulation [Sch79a]. Rejoinder [Key79b, Rob78b, Thi79b]. Relation [D’A72, Gue77e]. Relationship [Fue72, Pro73]. Relationships [Wei70]. Relative [EJO76]. Release [BR79]. Relevant [Dav76]. Remark [Bha73]. Remarks [CM76, Fed78, Gue78b]. Renowned [Ano71-42]. Report [Ano70b, Ano70c, Ano71c, Ano71-37, Ano72-39, Ano73z, Ano73b, Ano74j, Ano77k, CEF^{+}70, Eck72, Gol75, Kru71, Leh71, Leh72, Leh73, Mar76, Mos70, RWM^{+}72, Wei74, Ano76j, Ros70]. Reported [Haw74]. Reporting [HA75, WH71]. Representation [HM74, Hsi72, LL75, Wei70]. Required [Ano73p, Gok72]. Research [Ban71a, Ben72, Har74, BB71a]. Researchers [RJ73]. Resolution [Haw74]. Response [BPJ73, MH76]. Restriction [AM74, And75]. Restrictions [Man73b]. Result [Des71a]. Results [Bli71, HA75, Mor73, Kap79]. Reuben [Bac77]. Review [Car78a, Car78b, Fau77, Fau78, Gri79, GK79, IK78, KK78, Men79, Nam78, Nel79, Noe77, Pos78b, Pos78a, Pos79, Ros79, Shu78b, Shu78a, Sog79, SL79, Tra78, VK79]. Reviews [Pos74, Sea73]. Revisited [Bai77]. Rewards [Ban70]. Ridge [MS75]. Riefler [Gar74]. Right [JKS73]. Riley [Bar73]. Robert [Ano71-38]. Robust [Mal79]. Robustness [Ano79f, Pol78, Hog79]. Role [Bro74b, Ede79, Mos76, Ric77]. Ronald [McC75]. Root [KBG75]. Rootogram [Wai74]. Roots [MSH76]. Round [Shu78b, Shu78a]. Routine [Hua76]. Rubin [Mos79b]. Runs [Gue78b, Mar70].

S [Eis75]. S. [Eis75]. Sample [Ano73p, BPJ73, Bro74a, CS79, Eat78, FW76, GM78, Gok72, GK73b, Gue70, Gue75, Gue77b, Gue77d, Gue77c, Gue79, Lev79, Man75, Mar70, Max74, WD77, Tor78]. Sampling [Cut76, God70, HR71, Rob71, Str79, Wil73a]. Samuelson [WS79]. Sanders [Car78a]. Satellite [MN71]. Saved [Har74]. Scaled [JB70]. Scatters [BB71b]. Scheffe [CM76, Joh73, Joh74a]. Scheme [Gas77]. Scholarship [Bir71]. School [Pie76]. Schwarz [Arn74]. Science [Man74]. Sciences [Kru71, WC70]. Scientific [Bro76]. Scientist [Bro74b]. Seasonal [Men71, Rus76]. Seasonality [GD74]. Second [Ano70-39, Ano72-40, Ano73-39, Sei75]. Secretary [Ano70c, Leh71, Leh72, Leh73]. Section [CEF^{+}70]. Segmenting [Urq75]. Selection [GOS79, Rob71]. Sequence [PRPM79]. Sequential [Gro71]. Serial [Man75]. Series [BB71b, Fra77, Law79, Mee78, Nel76, RLM74, Sha79, Tra78]. Servants [Ben72]. Service [Ban71a, Ban72a, Urq75]. Set [Cra73, Gri79, Ryb75]. Sets [Fau78, Guj70b, Guj70a]. Several [Sei75]. Sex [Fue72]. Shapes [Bra77]. Sharing [AAF72, HS76b]. Sharpening [PW75]. Shiskin [Nor79]. Shoe [Bro74b]. Short [Bry72]. Should [BS75]. Showing [Cra73]. SIBYL [MW78]. SIBYL/RUNNER [MW78]. Sign [ES79, Gro71]. Significance [Blo78, Cra72]. Significant [And72]. Simeon [Mul77]. Simple [Blu75, GT71, Mar70, Mul72, ST73, Shu73, Sol75]. Simplified [Smi78]. Simplifying [SU72]. SIMSAM [Bri78]. Simulator [SL79]. Single [Rub73a, Sel71]. Situation [Bra77]. Six [BF78]. Six-State [BF78]. Size [Ano73p, CS79, Eat78, Fri72, Gok72, Gue70, Gue75, Gue77d, Gue77c, Gue79, Jan71, Wio72, Tor78]. Sleuthing [Wio75]. Smirnov [GM78, GK73b]. Smooth [RW74]. Snedecor [Ban74]. Social [Coc71, Kru71, Ros71, Tem73, WC70]. Societies [Ano72-28, Ano73-27]. Solution [SS70, Car78a]. Solving [BB75]. Some [Ano78e, Arn72, Bar74, Bha73, Bli71, CM76, Cro72, Cut76, Eck72, Fed78, FW76, Fol70, Fra77, Fre71, Gel78, Gin71, Gue77e, Gue78a, Gue78b, Hig75, Hun77, Kea74, Kot75, LB71, Lee78, Mal78, Mal79, MHF75, Mor73, Rao71, Rao73, Sas70, TB73, Whi73]. Special [Car78b, Fau77]. Specification [Gue71a, HMC78]. Specified [Ano73p, Gok72]. Speech [Ano71-40]. Splines [Smi79]. Spouses [Rub73a]. Spring [Ano71d, Ano73-38]. SPSS [BWSW79]. Square [Bli71, Gue77d, HS76b, JB70, KBG75]. Squared [D’A72, DH72]. Squares [Bha76, Bus73, KBG75, Rid77, SS70]. Stancliffe [Wil76]. Standard [BZ76, D’A70, Des71a, GT71, Lor75]. State [Ano71d, BF78, RM77, Gif70]. Stationary [Con77]. Statistic [Chi70, DN73, GK73b, Joh74a, KT75, Wac76]. Statistical [Ano70a, Ano70v, Ano70w, Ano70x, Ano70y, Ano70z, Ano71a, Ano71u, Ano71v, Ano71w, Ano71x, Ano72a, Ano72v, Ano72w, Ano72x, Ano72y, Ano72z, Ano72-28, Ano73a, Ano73x, Ano73y, Ano73-27, Ano73z, Ano76j, Ans73, Ban71a, Ban72b, Blo78, Boe72, BF78, Bry76, Car78b, Cle71, DuM79, Dun76, Fau78, FF73, Fed78, FHV75, Fra77, Gas75, Gol75, Hei79, Hog72, Hog79, IS79, KJ73, Kym70, Mal78, Man74, Mar79b, Men70, MF77, Mos76, Mos70, Nel72, Nor76, Pos76, Pos77b, Rob71, RGW70, RJ73, Rya79, Smi79, Thi79a, Thi79b, Wel79, WD77, Ros79]. Statistician [Ano71-42, Ban70, Ban74, Bro74b, Har72a, Ste73b, VC76, Ede79, Fau77]. Statisticians [Ano70d, BS75, CEF^{+}70, Eis70, Ken76, Rob78a, Rob78b, Wel78, Win78]. Statistics [AM74, And75, Ano72-39, Ano73z, Ban72a, Bar77, Beh77, Ben72, Bir71, Bly72, Bra73, BB75, Bro76, CCF^{+}77, CEF^{+}70, Dia75, Dun77, Eck72, Fau77, Fie79, Fol70, Fri77, Fue72, Gen77, GM78, Gif70, Goo72, Hau73, Hig75, HA75, KK78, KT75, Kru74, Lig75, Mar79a, Nov75, Orl74, Pad77, Pie76, Pos78a, RB74, Ric77, Sav70, Sch79a, She78, Shi70, Shi71b, Sid74, SB71, Tan79, Tem73, Urq71, WM72, Wal76, WH71, BR78a, LH72, Ano71-37, Ano76j, Mar74, Mar76, Men79, RWM^{+}72, Wei74, Gri79, IK78]. Status [RM77]. Stephan [Dem71]. Steps [Mos79a]. Stepwise [BA76]. Stevens [Gre71]. Stock [Gre71]. Strategies [Bro76, Dav76]. Strategy [MBG75]. Strengthening [Bir71, Shi70]. Strong [Ano78e, Gel78]. structural [Har76]. Student’s [Eis79, Hun77, PRPM79, RJ73, She78]. Studies [Cle71]. Study [Ano71-39, BL78, FT79, FF73, Wio75]. Subclass [Sea79]. Subject [Ano79a]. Subjective [Bly72]. Successes [Ste73a]. Sufficiency [Ano77e, Sal73, SS76, Sol72]. Sufficient [Joh74a, KT75, Wac76]. Sufficiently [BG76]. Suggestions [Sav70]. Suitable [FT79]. Sum [Len73]. Summary [BB72, Wei74]. Sums [DD72]. Supplement [Pos77b]. Supplementary [Eis75]. Support [WM73]. Suppressor [Day72]. Sure [CJ73]. Surface [Hig75, MH76]. Survey [Ano70p, Ano70o, Ano70-39, Ano70-40, Ano71b, Ano71o, Ano71-40, Ano71-41, Ano72b, Ano72o, Ano72n, Ano72-40, Ano72-41, Ano73r, Ano73q, Ano73-39, Ano73-40, BL78, God70, HR71, KJ73, Mor73, BB72, Kru71].

Survey-Sampling [God70]. Surveys [Ano74j]. Suspended [Wai74]. SWEEP [Goo79]. Symposia [MN71]. System [BR79, IS79, MHW74, Mal78, RJ73, WM72, WM73, Ano76j, Dun76, Mos79a]. Systematic [Man75]. Systems [AAF72, RGW70].

T [Ano71-31, Bar70, Mos72]. TAB [Fox79]. Table [Add74, Ano74e, BWSW79, Dro74, GM76, Gue72, Key70, KK74, Lee78, MH71, MH75, Shu78b, Shu78a, Wio75]. Tables [Fox79, Gue77e, Sne74]. Taeuber [Coa74]. Tails [Man76a, RW74]. Take [GK79]. Tax [Man76b]. Teachers [Mos70]. Teaching [AAF72, AM74, And75, Ban72a, Ben72, Boe72, Cut76, DuM79, Fau77, Fei71, Gen77, Hei79, Hun77, Joh74a, KS72, Lee78, LW79, Men79, PRPM79, Rya79, Shu73, Thi79a, Thi79b, VK79, WM72, Wel79, MH76]. Technique [HMC78, Joh74a]. Techniques [Hig75, Sch71]. Tennis [CC74]. Tenth [Nel77]. Terms [Ano73n, Dos73, Har72b, LMR75]. Test [Ano79h, Bea76, Car78a, ES79, Gro71, Gue75, Gue78b, Kym70, Lev79, Wil78b, Har76]. Testing [Bha76, D’A72, Guj70b, Guj70a, Hsi72, Kut74, SC71, Sol72, Sol75]. Tests [Cra72, D’A72, EF77, Fuc79, GK73a, GM76, Gre77, Gue72, Gue77a, Gue77d, Lor75, Sol75, SM79, Tau72]. Their [Gue77e, Mal79, PH77, RM77]. Theorem [Ano73o, Bar77, Fis73, Her72, Zer78, Ait71, HS76a, Mak79]. Theory [BP75, Hoy72, JK76, Nor76, Ryb75, Sas70, TK76]. Thereof [Ano77e, SS76]. Third [Ano70-40, Ano71-41, Ano72-41, Ano73-40, Shu78b]. Three [Nam78, Mon79]. three-factor [Mon79]. Ties [Cra73, ES79]. Time [AAF72, BB71b, Law79, Mee78, Nel76, Sha79]. Time-Series [BB71b]. Time-Sharing [AAF72]. Timer [Haw74]. Times [Haw74, Mor73]. TIMESERIES [Sha79]. today [Hog79, Goo72]. Tolerance [Ano73p, Gok72, Gue70]. Tool [AAF72, Smi79]. Total [Mul72, Str79]. Tournaments [Wio72]. Track [Bai77]. Tragedy [HR71]. Training [CEF^{+}70, Man74]. Transactional [BF78]. Transformation [FW76, HMC78, Mil78, SU72, Win79, YY75]. Transformations [Man76b, MSH76, Wat73]. transition [Eis79]. Transparency [Car78a, Fau78]. Transparently [Wat73]. Treasurer [Ano70c, Leh71, Leh72, Leh73]. Trends [KJ73, Ros71, Rus76]. Trial [Ede79, Fri72]. Trimmed [PH77]. Trimming [MHL77]. Truncated [Sel71]. TSERIES [Mee78]. Tucker [Cad72]. Tukey [Mil78]. Tutorial [Goo79]. Two [Adc73, CM76, Coo72, EF77, GM78, Gue77a, Guj70b, Hoc73, JB70, JK76, Key79a, Key79b, KK78, KA75, Max74, Mon79, Nel72, Nis79, Rob79, SC71, Sne74, Woo76, Bar76]. Two-dimensional [Mon79]. Two-Sample [GM78]. Two-Way [CM76, Hoc73, Sne74]. Type [Fei71, Fri72, GM76]. Typical [SS70].

U.S [Dun77]. U.S. [Dun76]. U.S.S.R. [Gif70]. Ubiquity [Kru74]. UMP [Ano79h, Wil78b]. Unbalanced [SH76]. Unbiased [Arn72, Gue71a, Gue71b, Gue72, Gue78a, GT71, Sti72]. Unbiasedness [Blu75]. unconditional [Kap79]. Unequal [Sea79]. Unequal-Subclass-Numbers [Sea79]. Uniformity [Mun77]. Universities [MF77]. University [Ano71-37, Ban71a, Cle71, Ano71d]. University-Wide [Ban71a]. untitled [Pos79, Sog79]. Unweighted [SM79]. Up-Dated [Gli70]. Urn [JK76]. Use [Bar77, Gue72, Gue78b, Gue79, Guj70b, Guj70a, Mal79, SC71, SH76, Bar76, Hog79, Kap79]. Used [WD77]. Useful [Des71a, PRPM79, Smi79]. User [Ano73z, Mee78]. User-Oriented [Mee78]. Uses [Key79a, Key79b, Nis79, Rob79]. Using [DuM79, Gen77, Hei79, HMC78, Kot75, Man75, MSH76, Nel77, Rya79, Thi79a, Thi79b, Wel79].

Validation [Wai74]. Value [Scl72]. Values [GP75, Rao76, YY75]. Vanishing [Dro74]. Variable [Beh70, Blo78, Day72, Men71, YY75]. Variables [Ano78e, DH72, Gel78, Gin71, GD74, Guj70b, Guj70a, Han71, HS76a, JB70, Men71, Rao73, Ryb75, Seb73, SU72, Bar76, KP79]. Variance [Add74, Ano74e, Arn72, Blu75, BP75, CM72, Gue71a, Gue72, Gue78a, Hol74, Hua76, Man75, Mar70, Rut71, ST73, Sea79, Mon79]. Variances [Gue77a, Max74, PH77, Rut71]. Variants [JK76]. Variate [Ban71b]. Variates [MSH76]. Variational [Bib73]. Variations [MTL78]. Various [GM76, RB74]. Vectors [SC71]. Versa [Ano73n, Dos73, Har72b]. Version [Ait71, Gli70]. versus [Sol75]. Vice [Ano73n, Dos73, Har72b]. Videotape [Nel79]. view [Hog79, LP76]. Viewpoint [Cut76]. Views [Mer71, Wei74]. vis [Ste73b]. vis-a-vis [Ste73b]. Visiting [Sid74]. Visual [Pos76, Wai74, Whi73]. Volume [Ano77l, Ano78l, Ano79o, Eck72]. Volumes [Ano79a]. vs [Ano79f, Bly72, Pol78].

W [Ban74, Gar74]. W. [Ano71-42]. Waiting [Mor73]. Walsh [Ano72-27]. Warren [You79]. Waste [Sal73]. Way [CM76, Hoc73, Sne74, Spe74]. Weak [Ano78e, Gel78]. Weighted [Rao77]. Which [Day72, LB71, Pad77]. Who [Har75, NJ73]. Whose [RW74]. Wide [Ban71a]. Widely [WD77]. Width [Des71b]. Wildlife [Kab70]. Wilks [Eis75]. William [Ano70-41]. Wind [Wio75]. Winfield [Gar74]. Winning [BR78b]. without [Ano73o, CJ73, Her72]. Witness [VC76]. Work [Ano72-39, Mar74]. Workers [Har74]. Workshop [Men79]. Workspace [Fox79].

Yarnold [Eat78]. Years [Rus76]. Youden [Ano71-42]. You’ve [CJ73].

Zero [Add74, Ano74e, Hun72, Kym70, MBG75, Wil73b, YY75]. Zero-Memory [MBG75]. Zyskind [Kem75].

References

Abranovic:1972:TSC

[AAF72] Wynn Abranovic, Roy Ageloff, and Donald Frederick. Time-sharing computer systems as a teaching tool. The American Statistician, 26(1):34–38, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682845.

Abramson:1973:MCM

[Abr73] Lee R. Abramson. Moments of a card matching distribution. The American Statistician, 27(4):166, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684049.

Adcock:1973:CTN

[Adc73] C. J. Adcock. Comparing two normal means. The American Statistician, 27(5):245, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087422.

Addelman:1974:CAV

[Add74] Sidney Addelman. Computing the analysis of variance table for experiments involving qualitative factors and zero amounts of quantitative factors. The American Statistician, 28(1):21–22, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683524.

Aitchison:1971:GVB

[Ait71] John Aitchison. A geometrical version of Bayes’ Theorem. The American Statistician, 25(5):45–46, December 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686084.

Anderson:1974:REA

[AM74] Virgil L. Anderson and Robert A. McLean. Restriction errors: Another dimension in teaching experimental statistics. The American Statistician, 28(4):145–152, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683342. See corrigendum [And75].

Anderson:1972:OCL

[And72] Donald A. Anderson. Overall confidence levels of the least significant difference procedure. The American Statistician, 26(4):30–32, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683204.

Anderson:1975:CRE

[And75] Virgil L. Anderson. Corrigendum: Restriction errors: Another dimension in teaching experimental statistics. The American Statistician, 29(3):134, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683444. See [AM74].

Anonymous:1970:AMJ

[Ano70a] Anonymous. 130th Annual Meeting: Joint Statistical Meetings. The American Statistician, 24(4):32–45, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681315.

Anonymous:1970:RBD

[Ano70b] Anonymous. 1969 report of the Board of Directors. The American Statistician, 24(2):43–44, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681937.

Anonymous:1970:RST

[Ano70c] Anonymous. 1969 report of the Secretary–Treasurer. The American Statistician, 24(2):44–46, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681938.

Anonymous:1970:ACA

[Ano70d] Anonymous. Addenda and corrigenda: Anniversaries of interest to statisticians. The American Statistician, 24(5):36, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682452.

Anonymous:1970:BMa

[Ano70e] Anonymous. Back matter. The American Statistician, 24(1):64, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682305.

Anonymous:1970:BMb

[Ano70f] Anonymous. Back matter. The American Statistician, 24(2):47–48, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681939.

Anonymous:1970:BMc

[Ano70g] Anonymous. Back matter. The American Statistician, 24(3):32, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682423.

Anonymous:1970:BMd

[Ano70h] Anonymous. Back matter. The American Statistician, 24(4):??, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681319.

Anonymous:1970:BMe

[Ano70i] Anonymous. Back matter. The American Statistician, 24(5):46–48, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682455.

Anonymous:1970:CNa

[Ano70j] Anonymous. Chapter notes. The American Statistician, 24(1):62– 63, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682304.

Anonymous:1970:CNb

[Ano70k] Anonymous. Chapter notes. The American Statistician, 24(2):41– 42, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681933.

Anonymous:1970:CNc

[Ano70l] Anonymous. Chapter notes. The American Statistician, 24(3):30– 31, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682422.

Anonymous:1970:CNd

[Ano70m] Anonymous. Chapter notes. The American Statistician, 24(4):55– 56, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681318.

Anonymous:1970:CNe

[Ano70n] Anonymous. Chapter notes. The American Statistician, 24(5): 44–45, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682454.

Anonymous:1970:FQSb

[Ano70o] Anonymous. First quarter survey of the economic outlook. The American Statistician, 24(2):40, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681932.

Anonymous:1970:FQSa

[Ano70p] Anonymous. Fourth quarter survey of the economic outlook. The American Statistician, 24(1):56–57, February 1970. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682302.

Anonymous:1970:FMa

[Ano70q] Anonymous. Front matter. The American Statistician, 24(1):10–13, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682286.

Anonymous:1970:FMb

[Ano70r] Anonymous. Front matter. The American Statistician, 24(2):1, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681919.

Anonymous:1970:FMc

[Ano70s] Anonymous. Front matter. The American Statistician, 24(3):??, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682412.

Anonymous:1970:FMd

[Ano70t] Anonymous. Front matter. The American Statistician, 24(4):1–8, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681303.

Anonymous:1970:FMe

[Ano70u] Anonymous. Front matter. The American Statistician, 24(5):1, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682438.

Anonymous:1970:ISAa

[Ano70v] Anonymous. International statistical activities. The American Statistician, 24(1):12–14, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682288.

Anonymous:1970:ISAb

[Ano70w] Anonymous. International statistical activities. The American Statistician, 24(2):8–9, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681922.

Anonymous:1970:ISAc

[Ano70x] Anonymous. International statistical activities. The American Statistician, 24(3):6–7, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682415.

Anonymous:1970:ISAd

[Ano70y] Anonymous. International statistical activities. The American Statistician, 24(4):9, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681306.

Anonymous:1970:ISAe

[Ano70z] Anonymous. International statistical activities. The American Statistician, 24(5):6–7, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682440.

Anonymous:1970:JCR

[Ano70-27] Anonymous. Jerome C. R. Li, 1914–1969. The American Statistician, 24(2):42, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681935.

Anonymous:1970:JBL

[Ano70-28] Anonymous. John B. Lansing, 1919–1970. The American Statistician, 24(5):17, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682445.

Anonymous:1970:Na

[Ano70-29] Anonymous. News. The American Statistician, 24(1):1–9, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682287.

Anonymous:1970:Nb

[Ano70-30] Anonymous. News. The American Statistician, 24(2):2–7, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681920.

Anonymous:1970:Nc

[Ano70-31] Anonymous. News. The American Statistician, 24(3):1–5, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682413.

Anonymous:1970:Nd

[Ano70-32] Anonymous. News. The American Statistician, 24(4):1–7, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681304.

Anonymous:1970:Ne

[Ano70-33] Anonymous. News. The American Statistician, 24(5):2–5, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682439.

Anonymous:1970:PNa

[Ano70-34] Anonymous. Personal news. The American Statistician, 24(1):14, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682289.

Anonymous:1970:PNb

[Ano70-35] Anonymous. Personal news. The American Statistician, 24(2):7, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681921.

Anonymous:1970:PNc

[Ano70-36] Anonymous. Personal news. The American Statistician, 24(3):5, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682414.

Anonymous:1970:PNd

[Ano70-37] Anonymous. Personal news. The American Statistician, 24(4):8, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2681305.

Anonymous:1970:PNe

[Ano70-38] Anonymous. Personal news. The American Statistician, 24(5):7–32, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682441.

Anonymous:1970:SQS

[Ano70-39] Anonymous. Second quarter survey of the economic outlook. The American Statistician, 24(3):7, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682416.

Anonymous:1970:TQS

[Ano70-40] Anonymous. Third quarter survey of the economic outlook. The American Statistician, 24(4):54, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681317.

Anonymous:1970:WF

[Ano70-41] Anonymous. William Feller, 1907–1970. The American Statistician, 24(2):42, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681934.

Anonymous:1971:AMJ

[Ano71a] Anonymous. 131st Annual Meeting: Joint Statistical Meetings. The American Statistician, 25(3):34–45, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683324.

Anonymous:1971:FQSa

[Ano71b] Anonymous. 1970 fourth quarter survey of the economic outlook. The American Statistician, 25(1):37, February 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682210.

Anonymous:1971:RBD

[Ano71c] Anonymous. 1970 report of the Board of Directors. The American Statistician, 25(3):57, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683329.

Anonymous:1971:SRM

[Ano71d] Anonymous. 1971 Spring Regional Meetings April 21–23 1971, Pennsylvania State University. The American Statistician, 25(1):6, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682202.

Anonymous:1971:BMa

[Ano71e] Anonymous. Back matter. The American Statistician, 25(1):??, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682223.

Anonymous:1971:BMb

[Ano71f] Anonymous. Back matter. The American Statistician, 25(2):46–48, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682010.

Anonymous:1971:BMc

[Ano71g] Anonymous. Back matter. The American Statistician, 25(3):64, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683332.

Anonymous:1971:BMd

[Ano71h] Anonymous. Back matter. The American Statistician, 25(4):??, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682933.

Anonymous:1971:BMe

[Ano71i] Anonymous. Back matter. The American Statistician, 25(5):??, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686091.

Anonymous:1971:CNa

[Ano71j] Anonymous. Chapter notes. The American Statistician, 25(1):64, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682222.

Anonymous:1971:CNb

[Ano71k] Anonymous. Chapter notes. The American Statistician, 25(2):44– 45, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682009.

Anonymous:1971:CNc

[Ano71l] Anonymous. Chapter notes. The American Statistician, 25(3):60– 63, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683331.

Anonymous:1971:CNd

[Ano71m] Anonymous. Chapter notes. The American Statistician, 25(4):54– 56, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682932.

Anonymous:1971:CNe

[Ano71n] Anonymous. Chapter notes. The American Statistician, 25(5):64, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686090.

Anonymous:1971:FQSb

[Ano71o] Anonymous. First quarter survey of the economic outlook. The American Statistician, 25(2):11, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681997.

Anonymous:1971:FMa

[Ano71p] Anonymous. Front matter. The American Statistician, 25(1):1–14, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682200.

Anonymous:1971:FMb

[Ano71q] Anonymous. Front matter. The American Statistician, 25(2):1–42, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681994.

Anonymous:1971:FMc

[Ano71r] Anonymous. Front matter. The American Statistician, 25(3):1–12, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683312.

Anonymous:1971:FMd

[Ano71s] Anonymous. Front matter. The American Statistician, 25(4):1–46, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682914.

Anonymous:1971:FMe

[Ano71t] Anonymous. Front matter. The American Statistician, 25(5):1–55, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686070.

Anonymous:1971:ISAa

[Ano71u] Anonymous. International statistical activities. The American Statistician, 25(1):11, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682203.

Anonymous:1971:ISAb

[Ano71v] Anonymous. International statistical activities. The American Statistician, 25(2):43, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682008.

Anonymous:1971:ISAc

[Ano71w] Anonymous. International statistical activities. The American Statistician, 25(3):9–10, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683314.

Anonymous:1971:ISAd

[Ano71x] Anonymous. International statistical activities. The American Statistician, 25(5):9, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686072.

Anonymous:1971:Na

[Ano71y] Anonymous. News. The American Statistician, 25(1):2–5, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682201.

Anonymous:1971:Nb

[Ano71z] Anonymous. News. The American Statistician, 25(2):2–8, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681995.

Anonymous:1971:Nc

[Ano71-27] Anonymous. News. The American Statistician, 25(3):2–8, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683313.

Anonymous:1971:Nd

[Ano71-28] Anonymous. News. The American Statistician, 25(4):2–8, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682915.

Anonymous:1971:Ne

[Ano71-29] Anonymous. News. The American Statistician, 25(5):3–8, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2686071.

Anonymous:1971:PM

[Ano71-30] Anonymous. Paul L. Meyer, 1926–1970. The American Statistician, 25(3):51, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683327.

Anonymous:1971:PB

[Ano71-31] Anonymous. Paul T. Bruyere. The American Statistician, 25(5):17, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686076.

Anonymous:1971:PNa

[Ano71-32] Anonymous. Personal news. The American Statistician, 25(1):11– 12, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682204.

Anonymous:1971:PNb

[Ano71-33] Anonymous. Personal news. The American Statistician, 25(2):8, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681996.

Anonymous:1971:PNc

[Ano71-34] Anonymous. Personal news. The American Statistician, 25(3):13, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683315.

Anonymous:1971:PNd

[Ano71-35] Anonymous. Personal news. The American Statistician, 25(4):8–9, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682916.

Anonymous:1971:PNe

[Ano71-36] Anonymous. Personal news. The American Statistician, 25(5):9–10, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686073.

Anonymous:1971:REC

[Ano71-37] Anonymous. Report of the Evaluation Committee on the University of Chicago Department of Statistics. The American Statistician, 25(3):17–24, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683317.

Anonymous:1971:RJ

[Ano71-38] Anonymous. Robert E. Johnson, 1905–1970. The American Statistician, 25(1):62, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682221.

Anonymous:1971:SFG

[Ano71-39] Anonymous. A study of future goals of ASA. The American Statistician, 25(4):11–16, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682918.

Anonymous:1971:SEO

[Ano71-40] Anonymous. Survey of the economic outlook after President Nixon’s speech on new economic policies. The American Statistician, 25 (5):11–12, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686074.

Anonymous:1971:TQS

[Ano71-41] Anonymous. Third quarter survey of the economic outlook. The American Statistician, 25(4):10, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682917.

Anonymous:1971:WJY

[Ano71-42] Anonymous. W. J. Youden, Renowned Statistician, 1907–1971. The American Statistician, 25(3):51, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683326.

Anonymous:1972:AMJ

[Ano72a] Anonymous. 132nd Annual Meeting: Joint Statistical Meetings. The American Statistician, 26(3):47–58, June 1972. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682869.

Anonymous:1972:ASE

[Ano72b] Anonymous. Annual survey of the economic outlook. The American Statistician, 26(4):11–12, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683199.

Anonymous:1972:BMa

[Ano72c] Anonymous. Back matter. The American Statistician, 26(1):??, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682853.

Anonymous:1972:BMb

[Ano72d] Anonymous. Back matter. The American Statistician, 26(2):??, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683465.

Anonymous:1972:BMc

[Ano72e] Anonymous. Back matter. The American Statistician, 26(3):??, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682874.

Anonymous:1972:BMd

[Ano72f] Anonymous. Back matter. The American Statistician, 26(4):??, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683216.

Anonymous:1972:BMe

[Ano72g] Anonymous. Back matter. The American Statistician, 26(5):??, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683790.

Anonymous:1972:CNa

[Ano72h] Anonymous. Chapter notes. The American Statistician, 26(1):55– 56, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682852.

Anonymous:1972:CNb

[Ano72i] Anonymous. Chapter notes. The American Statistician, 26(2):55– 56, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683464.

Anonymous:1972:CNc

[Ano72j] Anonymous. Chapter notes. The American Statistician, 26(3):71– 72, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682873.

Anonymous:1972:CNd

[Ano72k] Anonymous. Chapter notes. The American Statistician, 26(4):61– 64, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683215.

Anonymous:1972:CNe

[Ano72l] Anonymous. Chapter notes. The American Statistician, 26(5): 47–48, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683789.

Anonymous:1972:EC

[Ano72m] Anonymous. Editorial collaborators. The American Statistician, 26(5):36, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683785.

Anonymous:1972:FQSb

[Ano72n] Anonymous. First quarter survey of the economic outlook. The American Statistician, 26(2):10, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683450.

Anonymous:1972:FQSa

[Ano72o] Anonymous. Fourth quarter survey of the economic outlook. The American Statistician, 26(1):14, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682839.

Anonymous:1972:FMa

[Ano72p] Anonymous. Front matter. The American Statistician, 26(1):1–29, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682835.

Anonymous:1972:FMb

[Ano72q] Anonymous. Front matter. The American Statistician, 26(2):1–54, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683446.

Anonymous:1972:FMc

[Ano72r] Anonymous. Front matter. The American Statistician, 26(3):??, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682854.

Anonymous:1972:FMd

[Ano72s] Anonymous. Front matter. The American Statistician, 26(4):9–51, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683194.

Anonymous:1972:FMe

[Ano72t] Anonymous. Front matter. The American Statistician, 26(5):1–40, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683768.

Anonymous:1972:GN

[Ano72u] Anonymous. George E. Nicholson, 1918–1971. The American Statistician, 26(1):46, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682850.

Anonymous:1972:ISAa

[Ano72v] Anonymous. International statistical activities. The American Statistician, 26(1):12–13, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682838.

Anonymous:1972:ISAb

[Ano72w] Anonymous. International statistical activities. The American Statistician, 26(2):9, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683449.

Anonymous:1972:ISAc

[Ano72x] Anonymous. International statistical activities. The American Statistician, 26(3):6–7, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682857.

Anonymous:1972:ISAd

[Ano72y] Anonymous. International statistical activities. The American Statistician, 26(4):7–8, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683197.

Anonymous:1972:ISAe

[Ano72z] Anonymous. International statistical activities. The American Statistician, 26(5):12, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683771.

Anonymous:1972:JW

[Ano72-27] Anonymous. John E. Walsh, 1919–1972. The American Statistician, 26(5):41, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683787.

Anonymous:1972:MMC

[Ano72-28] Anonymous. Minutes of meetings of Committee of Presidents of Statistical Societies (COPSS). The American Statistician, 26(2):38– 42, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683462.

Anonymous:1972:Na

[Ano72-29] Anonymous. News. The American Statistician, 26(1):2–10, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682836.

Anonymous:1972:Nb

[Ano72-30] Anonymous. News. The American Statistician, 26(2):2–8, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683447.

Anonymous:1972:Nc

[Ano72-31] Anonymous. News. The American Statistician, 26(3):1–5, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682855.

Anonymous:1972:Nd

[Ano72-32] Anonymous. News. The American Statistician, 26(4):1–5, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683195.

Anonymous:1972:Ne

[Ano72-33] Anonymous. News. The American Statistician, 26(5):2–8, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683769.

Anonymous:1972:PNa

[Ano72-34] Anonymous. Personal news. The American Statistician, 26(1):10, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682837.

Anonymous:1972:PNb

[Ano72-35] Anonymous. Personal news. The American Statistician, 26(2):8, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683448.

Anonymous:1972:PNc

[Ano72-36] Anonymous. Personal news. The American Statistician, 26(3):5, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682856.

Anonymous:1972:PNd

[Ano72-37] Anonymous. Personal news. The American Statistician, 26(4):6, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683196.

Anonymous:1972:PNe

[Ano72-38] Anonymous. Personal news. The American Statistician, 26(5):8–11, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683770.

Anonymous:1972:PGW

[Ano72-39] Anonymous. Preparation for graduate work in statistics. A report of the CUPM panel on statistics. The American Statistician, 26(1):19–25, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682841.

Anonymous:1972:SQS

[Ano72-40] Anonymous. Second quarter 1972 survey of the economic outlook. The American Statistician, 26(4):10–11, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683198.

Anonymous:1972:TQS

[Ano72-41] Anonymous. Third quarter 1972 survey of the economic outlook. The American Statistician, 26(5):13, December 1972. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683772.

Anonymous:1973:AMJ

[Ano73a] Anonymous. 133rd Annual Meeting: Joint Statistical Meetings. The American Statistician, 27(4):172–185, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684053.

Anonymous:1973:RBD

[Ano73b] Anonymous. 1972 report of the Board of Directors. The American Statistician, 27(3):128–129, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683641.

Anonymous:1973:BMa

[Ano73c] Anonymous. Back matter. The American Statistician, 27(1):46–47, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682913.

Anonymous:1973:BMb

[Ano73d] Anonymous. Back matter. The American Statistician, 27(2):??, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683971.

Anonymous:1973:BMc

[Ano73e] Anonymous. Back matter. The American Statistician, 27(3):??, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683644.

Anonymous:1973:BMd

[Ano73f] Anonymous. Back matter. The American Statistician, 27(4):200, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684057.

Anonymous:1973:BMe

[Ano73g] Anonymous. Back matter. The American Statistician, 27(5):1–248, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/3087427.

Anonymous:1973:CNa

[Ano73h] Anonymous. Chapter notes. The American Statistician, 27(1):45– 48, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682912.

Anonymous:1973:CNb

[Ano73i] Anonymous. Chapter notes. The American Statistician, 27(2):96, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683970.

Anonymous:1973:CNc

[Ano73j] Anonymous. Chapter notes. The American Statistician, 27(3):134– 136, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683643.

Anonymous:1973:CNd

[Ano73k] Anonymous. Chapter notes. The American Statistician, 27(4): 196–199, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684056.

Anonymous:1973:CNe

[Ano73l] Anonymous. Chapter notes. The American Statistician, 27(5):247, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/3087426.

Anonymous:1973:CLE

[Ano73m] Anonymous. Corrections: Letters to the Editor. The American Statistician, 27(1):44, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682909.

Anonymous:1973:CEM

[Ano73n] Anonymous. Corrections: “On Expressing Moments in Terms of Cumulants and Vice Versa”. The American Statistician, 27(1):44, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682910.

See [Har72b].

Anonymous:1973:CPP

[Ano73o] Anonymous. Corrections: Properties of the Pascal distribution without recourse to the binomial theorem with negative exponent.

The American Statistician, 27(1):44, February 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682911. See [Her72].

Anonymous:1973:CFS

[Ano73p] Anonymous. Corrigendum to a formula for sample size required for specified tolerance. The American Statistician, 27(3):127, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683640.

Anonymous:1973:FQSb

[Ano73q] Anonymous. First quarter 1973 survey of the economic outlook. The American Statistician, 27(2):57, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683957.

Anonymous:1973:FQSa

[Ano73r] Anonymous. Fourth quarter 1972 survey of the economic outlook. The American Statistician, 27(1):11, February 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682896.

Anonymous:1973:FMa

[Ano73s] Anonymous. Front matter. The American Statistician, 27(1):1–38, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682892.

Anonymous:1973:FMb

[Ano73t] Anonymous. Front matter. The American Statistician, 27(2):56–95, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683953.

Anonymous:1973:FMc

[Ano73u] Anonymous. Front matter. The American Statistician, 27(3):??, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683626.

Anonymous:1973:FMd

[Ano73v] Anonymous. Front matter. The American Statistician, 27(4):137, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2684038.

Anonymous:1973:FMe

[Ano73w] Anonymous. Front matter. The American Statistician, 27(5): 201–208, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087398.

Anonymous:1973:ISAa

[Ano73x] Anonymous. International statistical activities. The American Statistician, 27(1):9–10, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682895.

Anonymous:1973:ISAb

[Ano73y] Anonymous. International statistical activities. The American Statistician, 27(3):101–102, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683629.

Anonymous:1973:MPI

[Ano73z] Anonymous. Maintaining the professional integrity of federal statistics. A report of the American Statistical Association–Federal Statistics Users’ Conference Committee on the integrity of federal statistics. The American Statistician, 27(2):58–67, April 1973. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683958.

Anonymous:1973:MCP

[Ano73-27] Anonymous. Minutes of Committee of Presidents of Statistical Societies. The American Statistician, 27(2):88–90, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683968.

Anonymous:1973:Na

[Ano73-28] Anonymous. News. The American Statistician, 27(1):2–7, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682893.

Anonymous:1973:Nb

[Ano73-29] Anonymous. News. The American Statistician, 27(2):49–52, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683954.

Anonymous:1973:Nc

[Ano73-30] Anonymous. News. The American Statistician, 27(3):97–100, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683627.

Anonymous:1973:Nd

[Ano73-31] Anonymous. News. The American Statistician, 27(4):138–143, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684039.

Anonymous:1973:Ne

[Ano73-32] Anonymous. News. The American Statistician, 27(5):202–206, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/3087399.

Anonymous:1973:PNa

[Ano73-33] Anonymous. Personal news. The American Statistician, 27(1):8, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682894.

Anonymous:1973:PNb

[Ano73-34] Anonymous. Personal news. The American Statistician, 27(2):53, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683955.

Anonymous:1973:PNc

[Ano73-35] Anonymous. Personal news. The American Statistician, 27(3):101, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683628.

Anonymous:1973:PNd

[Ano73-36] Anonymous. Personal news. The American Statistician, 27(4): 143–144, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684040.

Anonymous:1973:PNe

[Ano73-37] Anonymous. Personal news. The American Statistician, 27(5): 206–209, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087400.

Anonymous:1973:PPS

[Ano73-38] Anonymous. Preliminary program: 1973 Spring Regional Meetings. The American Statistician, 27(2):54–55, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683956.

Anonymous:1973:SQS

[Ano73-39] Anonymous. Second quarter 1973 survey of the economic outlook. The American Statistician, 27(4):146–147, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684042.

Anonymous:1973:TQS

[Ano73-40] Anonymous. Third quarter 1973 survey of the economic outlook. The American Statistician, 27(5):210–211, December 1973. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/3087401.

Anonymous:1974:BMa

[Ano74a] Anonymous. Back matter. The American Statistician, 28(1):??, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683531.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. The American Statistician, 28(2):78–88, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683610.

Anonymous:1974:BMc

[Ano74c] Anonymous. Back matter. The American Statistician, 28(3):??, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683043.

Anonymous:1974:BMd

[Ano74d] Anonymous. Back matter. The American Statistician, 28(4): 167–176, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683347.

Anonymous:1974:CCA

[Ano74e] Anonymous. Correction: Computing the analysis of variance table for experiments involving qualitative factors and zero amounts of quantitative factors. The American Statistician, 28(3):??, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683042.

Anonymous:1974:FMa

[Ano74f] Anonymous. Front matter. The American Statistician, 28(1):1–26, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683517.

Anonymous:1974:FMb

[Ano74g] Anonymous. Front matter. The American Statistician, 28(2):41–55, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683596.

Anonymous:1974:FMc

[Ano74h] Anonymous. Front matter. The American Statistician, 28(3):??, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683030.

Anonymous:1974:FMd

[Ano74i] Anonymous. Front matter. The American Statistician, 28(4): 113–125, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683333.

Anonymous:1974:RAC

[Ano74j] Anonymous. Report on the ASA Conference on Surveys of Human Populations. The American Statistician, 28(1):30–34, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683529.

Anonymous:1975:BMa

[Ano75a] Anonymous. Back matter. The American Statistician, 29(1):72, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683690.

Anonymous:1975:BMb

[Ano75b] Anonymous. Back matter. The American Statistician, 29(2):111– 112, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683282.

Anonymous:1975:BMc

[Ano75c] Anonymous. Back matter. The American Statistician, 29(3):135– 136, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683445.

Anonymous:1975:BMd

[Ano75d] Anonymous. Back matter. The American Statistician, 29(4): 176–184, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683491.

Anonymous:1975:CBP

[Ano75e] Anonymous. Corrections: Blackjack probabilities. The American Statistician, 29(4):175, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683490. See [MBG75].

Anonymous:1975:FMa

[Ano75f] Anonymous. Front matter. The American Statistician, 29(1):1–2, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683672.

Anonymous:1975:FMb

[Ano75g] Anonymous. Front matter. The American Statistician, 29(2):??, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683267.

Anonymous:1975:FMc

[Ano75h] Anonymous. Front matter. The American Statistician, 29(3):??, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683432.

Anonymous:1975:FMd

[Ano75i] Anonymous. Front matter. The American Statistician, 29(4):137, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683479.

Anonymous:1976:BMa

[Ano76a] Anonymous. Back matter. The American Statistician, 30(1):46–48, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682891.

Anonymous:1976:BMb

[Ano76b] Anonymous. Back matter. The American Statistician, 30(2):??, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683806.

Anonymous:1976:BMc

[Ano76c] Anonymous. Back matter. The American Statistician, 30(3):??, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683866.

Anonymous:1976:BMd

[Ano76d] Anonymous. Back matter. The American Statistician, 30(4): 207–216, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683767.

Anonymous:1976:CNR

[Ano76e] Anonymous. Correction: Nonlinear regression. The American Statistician, 30(1):44–45, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682890.

Anonymous:1976:FMa

[Ano76f] Anonymous. Front matter. The American Statistician, 30(1):1–39, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682877.

Anonymous:1976:FMb

[Ano76g] Anonymous. Front matter. The American Statistician, 30(2):76–77, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683791.

Anonymous:1976:FMc

[Ano76h] Anonymous. Front matter. The American Statistician, 30(3):??, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683853.

Anonymous:1976:FMd

[Ano76i] Anonymous. Front matter. The American Statistician, 30(4):153, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683750.

Anonymous:1976:IFS

[Ano76j] Anonymous. Improving the Federal Statistical System. Report of the Joint Ad Hoc Committee on Government Statistics. The American Statistician, 30(4):154–158, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683751.

Anonymous:1977:BMa

[Ano77a] Anonymous. Back matter. The American Statistician, 31(1):55–56, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682984.

Anonymous:1977:BMb

[Ano77b] Anonymous. Back matter. The American Statistician, 31(2):100, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683055.

Anonymous:1977:BMc

[Ano77c] Anonymous. Back matter. The American Statistician, 31(3):136– 140, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682967.

Anonymous:1977:BMd

[Ano77d] Anonymous. Back matter. The American Statistician, 31(4):198, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683547.

Anonymous:1977:CSM

[Ano77e] Anonymous. Corrigenda: Sufficiency, minimal sufficiency, and the lack thereof. The American Statistician, 31(1):54, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682983. See [SS76].

Anonymous:1977:EC

[Ano77f] Anonymous. Editorial collaborators. The American Statistician, 31(4):196–197, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683546.

Anonymous:1977:FMa

[Ano77g] Anonymous. Front matter. The American Statistician, 31(1):1–49, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682968.

Anonymous:1977:FMb

[Ano77h] Anonymous. Front matter. The American Statistician, 31(2):57–58, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683044.

Anonymous:1977:FMc

[Ano77i] Anonymous. Front matter. The American Statistician, 31(3):??, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682953.

Anonymous:1977:FMd

[Ano77j] Anonymous. Front matter. The American Statistician, 31(4): 184–194, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683533.

Anonymous:1977:RAH

[Ano77k] Anonymous. Report of Ad Hoc Committee on Privacy and Confidentiality. The American Statistician, 31(2):59–78, May 1977. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683045.

Anonymous:1977:VI

[Ano77l] Anonymous. Volume information. The American Statistician, 31(4):195–196, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683532.

Anonymous:1978:BMa

[Ano78a] Anonymous. Back matter. The American Statistician, 32(1):40–44, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683478.

Anonymous:1978:BMb

[Ano78b] Anonymous. Back matter. The American Statistician, 32(2):78–80, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683625.

Anonymous:1978:BMc

[Ano78c] Anonymous. Back matter. The American Statistician, 32(3):115– 116, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683363.

Anonymous:1978:BMd

[Ano78d] Anonymous. Back matter. The American Statistician, 32(4): 159–160, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682952.

Anonymous:1978:CSE

[Ano78e] Anonymous. Corrigenda: Some examples of the weak and strong laws of large numbers for averages of mutually independent random variables. The American Statistician, 32(4):146, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682950. See [Gel78].

Anonymous:1978:CMR

[Ano78f] Anonymous. Corrigenda: The hat matrix in regression and ANOVA. The American Statistician, 32(4):146, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682949. See [HW78].

Anonymous:1978:EC

[Ano78g] Anonymous. Editorial collaborators. The American Statistician, 32(4):157–158, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682951.

Anonymous:1978:FMa

[Ano78h] Anonymous. Front matter. The American Statistician, 32(1):??, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683466.

Anonymous:1978:FMb

[Ano78i] Anonymous. Front matter. The American Statistician, 32(2):??, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683611.

Anonymous:1978:FMc

[Ano78j] Anonymous. Front matter. The American Statistician, 32(3):??, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683348.

Anonymous:1978:FMd

[Ano78k] Anonymous. Front matter. The American Statistician, 32(4): 147–155, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682935.

Anonymous:1978:VI

[Ano78l] Anonymous. Volume information. The American Statistician, 32(4):156–157, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682934.

Anonymous:1979:ASI

[Ano79a] Anonymous. Author and subject index to volumes 1–32, 1947– 1978. The American Statistician, 33(5):249–308, December 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682876.

Anonymous:1979:BMa

[Ano79b] Anonymous. Back matter. The American Statistician, 33(1):38–39, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683075.

Anonymous:1979:BMb

[Ano79c] Anonymous. Back matter. The American Statistician, 33(2):95–96, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683234.

Anonymous:1979:BMc

[Ano79d] Anonymous. Back matter. The American Statistician, 33(3):??, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683829.

Anonymous:1979:BMd

[Ano79e] Anonymous. Back matter. The American Statistician, 33(4): 243–248, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683749.

Anonymous:1979:CIR

[Ano79f] Anonymous. Corrigenda: Inference robustness vs. criterion robustness: An example. The American Statistician, 33(1):37, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683074. See

[Pol78].

Anonymous:1979:CNA

[Ano79g] Anonymous. Corrigenda: On the non-attainability of Chebyshev bounds. The American Statistician, 33(1):37, February 1979. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683072. See [GM77a].

Anonymous:1979:CWC

[Ano79h] Anonymous. Corrigenda: When is the Chow test UMP? The American Statistician, 33(1):37, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683073. See [Wil78b].

Anonymous:1979:EC

[Ano79i] Anonymous. Editorial collaborators. The American Statistician, 33(4):240–242, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683748.

Anonymous:1979:FMa

[Ano79j] Anonymous. Front matter. The American Statistician, 33(1):??, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683056.

Anonymous:1979:FMb

[Ano79k] Anonymous. Front matter. The American Statistician, 33(2):??, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683217.

Anonymous:1979:FMc

[Ano79l] Anonymous. Front matter. The American Statistician, 33(3):??, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683807.

Anonymous:1979:FMd

[Ano79m] Anonymous. Front matter. The American Statistician, 33(4): 230–238, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683728.

Anonymous:1979:FMe

[Ano79n] Anonymous. Front matter. The American Statistician, 33(5):??, December 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682875.

Anonymous:1979:VI

[Ano79o] Anonymous. Volume information. The American Statistician, 33(4):239–240, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683727.

Anscombe:1973:GSA

[Ans73] F. J. Anscombe. Graphs in statistical analysis. The American Statistician, 27(1):17–21, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682899.

Arnold:1972:SEM

[Arn72] Barry C. Arnold. Some examples of minimum variance unbiased estimates. The American Statistician, 26(4):34–36, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683206.

Arnold:1974:SRE

[Arn74] Barry C. Arnold. Schwarz, regression, and extreme deviance. The American Statistician, 28(1):22–23, February 1974. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683525.

Bendel:1976:CSR

[BA76] R. B. Bendel and A. A. Afifi. A criterion for stepwise regression. The American Statistician, 30(2):85–87, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683800.

Backman:1977:MRG

[Bac77] Jules Backman. Martin Reuben Gainsbrugh, 1907–1977. The American Statistician, 31(4):178, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683543.

Bailey:1977:RTR

[Bai77] B. J. R. Bailey. The racing track revisited. The American Statistician, 31(1):30–33, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682975.

Bancroft:1970:WRM

[Ban70] T. A. Bancroft. What rewards may a statistician expect? The American Statistician, 24(5):8–10, December 1970. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682442.

Bancroft:1971:EUW

[Ban71a] T. A. Bancroft. On establishing a university-wide statistical consulting and cooperative research service. The American Statistician, 25(5):21–24, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686078.

Banerjee:1971:MCD

[Ban71b] K. S. Banerjee. On the marginal and conditional distributions of the multi-variate normal. The American Statistician, 25(5):50–51, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686086.

Bancroft:1972:TSC

[Ban72a] T. A. Bancroft. On teaching of service courses in statistics. The American Statistician, 26(2):14–16, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683452.

Bancroft:1972:SCP

[Ban72b] T. A. Bancroft. The statistical community and the protection of privacy. The American Statistician, 26(4):13–16, October 1972. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683200.

Bancroft:1974:GWS

[Ban74] T. A. Bancroft. George W. Snedecor, Pioneer Statistician, 1881– 1974. The American Statistician, 28(3):108–109, August 1974. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683039.

Barbour:1970:EC

[Bar70] Dana M. Barbour. Edward T. Crowder, 1910–1970. The American Statistician, 24(2):42, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681936.

Barbour:1973:DCR

[Bar73] Dana M. Barbour. Donald C. Riley, 1905–1973. The American Statistician, 27(4):145–146, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684041.

Barrett:1974:CDS

[Bar74] James P. Barrett. The coefficient of determination — some limitations. The American Statistician, 28(1):19–20, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683523.

Barnow:1976:UPV

[Bar76] Burt S. Barnow. The use of proxy variables when one or two independent variables are measured with error. The American Statistician, 30(3):119–121, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683856.

Bartels:1977:ULT

[Bar77] R. Bartels. On the use of limit theorem arguments in economic statistics. The American Statistician, 31(2):85–87, May 1977. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683048.

Berg:1971:ICN

[BB71a] Sanford Berg and Neville Beharie. Innovations in communication: The National Bureau of Economic Research and the computer. The American Statistician, 25(5):28–32, December 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686080.

Bry:1971:IAT

[BB71b] Gerhard Bry and Charlotte Boschan. Interpretation and analysis of time-series scatters. The American Statistician, 25(2):29–33, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682005.

Boardman:1972:APS

[BB72] Thomas J. Boardman and Eileen C. Boardman. ASA 1971 Publications Survey — summary of findings. The American Statistician, 26(4):17–22, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683201.

Brightman:1975:PSM

[BB75] Harvey Brightman and Michael Broida. On problem solving, motivation and statistics. The American Statistician, 29(4):164–166, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683485.

Berry:1972:LE

[BDR^{+}72] Donald A. Berry, W. Edwards Deming, A. C. Rosander, Robert A. Agnew, C. Philip Cox, L. R. LaMotte, Warren Dent, I. J. Good, Franklin H. Silverman, Hans K. Ury, Marian B. Ury, John Havas, S. R. Searle, D. S. Robson, Lawrence L. Kupper, Stanley W. Nash, A. L. Wilson, S. L. George, M. M. Desu, Ralph E. Beals, and Herbert A. David. Letters to the Editor. The American Statistician, 26(1):47–54, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682851.

Beal:1976:FHT

[Bea76] Stuart Beal. Fisher’s hypergeometric test for a comparison in a finite population. The American Statistician, 30(4):165–168, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683753.

Behboodian:1970:ERV

[Beh70] Javad Behboodian. On the expectation of the random variable (X^{2} − Y^{2})^{1/2}. The American Statistician, 24(4):28–29, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681312.

Behboodian:1977:NFF

[Beh77] Javad Behboodian. A note on functional forms of order statistics. The American Statistician, 31(4):161–162, November 1977. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683539.

Benjamin:1972:TSO

[Ben72] Bernard Benjamin. Teaching statistics and operational research to civil servants. The American Statistician, 26(4):23–26, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683202.

Betteley:1977:ALE

[Bet77] I. G. Betteley. The addition law for expectations. The American Statistician, 31(1):33–35, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682976.

Boen:1978:SST

[BF78] James Boen and David Fryd. Six-state transactional analysis in statistical consultation. The American Statistician, 32(2):58–60, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683616.

Barrett:1976:WSL

[BG76] James P. Barrett and Leland Goldsmith. When is n sufficiently large? The American Statistician, 30(2):67–70, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683795.

Bhapkar:1973:RIS

[Bha73] V. P. Bhapkar. A remark on the invariance of some estimators. The American Statistician, 27(5):231–232, December 1973. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/3087409.

Bhapkar:1976:GLS

[Bha76] V. P. Bhapkar. Generalized least squares estimation and testing. The American Statistician, 30(2):73–74, May 1976. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683797.

Bibby:1973:VPM

[Bib73] John Bibby. Variational proofs in multivariate analysis. The American Statistician, 27(5):244–245, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/3087421.

Birnbaum:1971:PSS

[Bir71] Allan Birnbaum. A perspective for strengthening scholarship in statistics. The American Statistician, 25(3):14–17, June 1971. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683316.

Bailar:1978:PSD

[BL78] Barbara A. Bailar and C. Michael Lanphier. A pilot study to develop survey methods to assess survey practices. The American Statistician, 32(4):130–132, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682939.

Blight:1971:SGR

[Bli71] B. J. N. Blight. Some general results on reduced mean square error estimation. The American Statistician, 25(3):24–25, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683318.

Blom:1976:WAM

[Blo76] Gunnar Blom. When is the arithmetic mean blue? The American Statistician, 30(1):40–42, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682888.

Bloch:1978:MES

[Blo78] Farrell E. Bloch. Measurement error and statistical significance of an independent variable. The American Statistician, 32(1):26–27, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683471.

Blumenthal:1975:SPL

[Blu75] S. Blumenthal. A simple proof of local minimum variance unbiasedness. The American Statistician, 29(3):130, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683441.

Blyth:1972:SVO

[Bly72] Colin R. Blyth. Subjective vs. objective methods in statistics. The American Statistician, 26(3):20–22, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682860.

Barnard:1972:LE

[BMF^{+}72] G. A. Barnard, Geoffrey H. Moore, James W. Frane, Campbell B. Read, Lawrence M. Aleamoni, Marvin M. Engel, W. Edwards Deming, Richard E. Cavicchi, Seymour Geisser, Hugh A. Reeder, James Koziol, David P. Byar, Manus Rabinowitz, Mark Berenson, David M. Grether, A. M. Kshirsagar, Elliot M. Cramer, Stanley L. Sclove, Robert I. Berg, Mitchell Snyder, and Dinesh S. Bhoj. Letters to the Editor. The American Statistician, 26(4):52–60, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683214.

Boen:1972:TPI

[Boe72] James R. Boen. The teaching of personal interaction in statistical consulting. The American Statistician, 26(1):30–31, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682843.

Bookstein:1975:FPL

[Boo75] Fred L. Bookstein. On a form of piecewise linear regression. The American Statistician, 29(3):116–117, August 1975. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683435.

Bohrer:1975:EAV

[BP75] Robert Bohrer and Richard F. Potthoff. An exercise in analysis of variance theory. The American Statistician, 29(4):166–168, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683486.

Brand:1973:LSC

[BPJ73] Richard J. Brand, Dudley E. Pinnock, and Kirby L. Jackson. Large sample confidence bands for the logistic response curve and its inverse. The American Statistician, 27(4):157–160, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684046.

Beniger:1978:QGS

[BR78a] James R. Beniger and Dorothy L. Robyn. Quantitative graphics in statistics: a brief history. The American Statistician, 32(1):1–11, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683467.

Berry:1978:PWC

[BR78b] Donald A. Berry and Ronald R. Regal. Probabilities of winning a certain carnival game. The American Statistician, 32(4):126–129, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682938.

Baker:1979:GSR

[BR79] R. J. Baker and M. G. Richardson. The GLIM System — Release 3. The American Statistician, 33(4):221–222, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683741.

Brager:1973:EQS

[Bra73] Gary L. Brager. Ethical questions in statistics. The American Statistician, 27(5):246, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/3087424.

Bradley:1977:CSC

[Bra77] James V. Bradley. A common situation conducive to bizarre distribution shapes. The American Statistician, 31(4):147–150, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683535.

Bridges:1978:S

[Bri78] George Bridges. SIMSAM. The American Statistician, 32(3):110, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683360.

Brown:1971:CCD

[Bro71] Robert W. C. Brown. Comparing census data in 90 countries. The American Statistician, 25(1):32–37, February 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682209.

Broffitt:1974:ELS

[Bro74a] James D. Broffitt. An example of the large sample behavior of the midrange. The American Statistician, 28(2):69–70, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683607.

Bross:1974:RSS

[Bro74b] Irwin D. J. Bross. The role of the statistician: Scientist or shoe clerk. The American Statistician, 28(4):126–127, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683335.

Bross:1976:SSH

[Bro76] Irwin D. J. Bross. Scientific strategies in human affairs: Explaining statistics to the public. The American Statistician, 30(4):171–175, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683755.

Bryant:1972:SPK

[Bry72] Peter Bryant. A short proof of a known lemma. The American Statistician, 26(5):35–36, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683783.

Bryson:1973:CCD

[Bry73] Maurice C. Bryson. Craps with crooked dice. The American Statistician, 27(4):167–168, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684050.

Bryson:1976:LDP

[Bry76] Maurice C. Bryson. The literary digest poll: Making of a statistical myth. The American Statistician, 30(4):184–185, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683758.

Boen:1975:SSC

[BS75] J. Boen and H. Smith. Should statisticians be certified? The American Statistician, 29(3):113–114, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683433.

Bradley:1979:CPR

[BS79] Ralph A. Bradley and Sushil S. Srivastava. Correlation in polynomial regression. The American Statistician, 33(1):11–14, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683059.

Buse:1973:GFG

[Bus73] A. Buse. Goodness of fit in generalized least squares estimation. The American Statistician, 27(3):106–108, June 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683631.

Brown:1979:LTA

[BWSW79] Barry W. Brown, Herman Walker, Marcus Schimek, and Peggy R. Wright. A life table analysis package for SPSS. The American Statistician, 33(4):225–227, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683746.

Bondy:1976:SEM

[BZ76] Warren H. Bondy and William Zlot. The standard error of the mean and the difference between means for finite populations. The American Statistician, 30(2):96–97, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683803.

Cady:1972:HT

[Cad72] Foster Cady. Henry Tucker, 1923–1971. The American Statistician, 26(1):46, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682849.

Cardenas:1978:BRB

[Car78a] Manuel Cardenas. Book review: Sanders, Murph, and Eng, Solution Manual, Test Questions and Transparency Masters. The American Statistician, 32(3):108, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683357.

Carlson:1978:RSI

[Car78b] Roger Carlson. Review of special issue on statistical education. The American Statistician, 32(4):139–140, November 1978. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682943.

Carter:1974:AGT

[CC74] Walter H. Carter, Jr. and Sharon L. Crews. An analysis of the game of tennis. The American Statistician, 28(4):130–134, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683337.

Chervany:1977:FDM

[CCF^{+}77] Norman L. Chervany, Raymond O. Collier, Jr., Stephen E. Fienberg, Paul E. Johnson, and John Neter. A framework for the development of measurement instruments for evaluating the introductory statistics course. The American Statistician, 31(1):17–23, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682972.

Colby:1970:ETS

[CEF^{+}70] Marian M. Colby, Carl L. Erhardt, Todd M. Frazier, Felix E. Moore, Carroll E. Palmer, William F. Taylor, and Iwao M. Moriyama. Education and training of statisticians for health agencies: a condensed report of the Committee of the Statistics Section, American Public Health Association. The American Statistician, 24(5):22–27, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682447.

Cicchetti:1979:RRA

[CH79] Domenic V. Cicchetti and Robert Heavens, Jr. RATCAT (rater agreement/categorical data). The American Statistician, 33(2):91, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683231.

Chew:1971:PEP

[Che71] Victor Chew. Point estimation of the parameter of the binomial distribution. The American Statistician, 25(5):47–50, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686085.

Chissom:1970:IKS

[Chi70] Brad S. Chissom. Interpretation of the kurtosis statistic. The American Statistician, 24(4):19–22, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681309.

Campbell:1973:HGA

[CJ73] Cathy Campbell and Brian L. Joiner. How to get the answer without being sure you’ve asked the question. The American Statistician, 27(5):229–231, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087408.

Clelland:1971:SSU

[Cle71] Richard C. Clelland. On statistical studies of university education. The American Statistician, 25(4):20–24, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682920.

Carter:1972:OCG

[CM72] Walter H. Carter, Jr. and Raymond H. Myers. Orthogonal contrasts and the generalized inverse in fixed effects analysis of variance. The American Statistician, 26(5):32–34, December 1972. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683781.

Cohen:1976:SRS

[CM76] Arthur Cohen and John J. Miller. Some remarks on Scheffe’s twoway mixed model. The American Statistician, 30(1):36–37, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682886.

Cohen:1978:LE

[CN78] Joel E. Cohen and John Neter. Letters to the Editor. The American Statistician, 32(1):39, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683477.

Crocker:1972:LE

[CNR^{+}72] Douglas C. Crocker, Subhash Narula, John S. Ramberg, Y. Horiba, Arlo Hendrickson, Elliot M. Cramer, T. M. F. Smith, Haynes Phillips, Raymond A. Gibley, Richard Sundstrom, Nathan Mantel, B. J. N. Blight, Johannes Gordesch, J. C. Ahuja, Colin White, Donald E. Henschel, A. P. Dempster, and Lawrence Harris. Letters to the Editor. The American Statistician, 26(5):41–47, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683788.

Coale:1974:IBT

[Coa74] Ansley J. Coale. Irene Barnes Taeuber, 1906–1974. The American Statistician, 28(3):109–110, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683040.

Cochran:1971:DRS

[Coc71] Robert S. Cochran. Designs with redeeming social aspects for evenings of social bridge. The American Statistician, 25(2):12–15, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681998.

Connell:1977:CMC

[Con77] Fred J. Connell. Construction of a Markov chain with given stationary distribution. The American Statistician, 31(2):93, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683051.

Cooper:1972:TAP

[Coo72] J. Phillip Cooper. Two approaches to polynomial distributed lags estimation: An expository note and comment. The American Statistician, 26(3):32–35, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682864.

Cowden:1975:MEC

[Cow75] Dudley J. Cowden. A method of evaluating contestants. The American Statistician, 29(2):82–84, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683269.

Cramer:1972:STT

[Cra72] Elliot M. Cramer. Significance tests and tests of models in multiple regression. The American Statistician, 26(4):26–30, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683203.

Crain:1973:NSL

[Cra73] Bradford R. Crain. A note on showing a lack of ties in a set of s observations. The American Statistician, 27(4):169, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684051.

Crocker:1972:SIM

[Cro72] Douglas C. Crocker. Some interpretations of the multiple correlation coefficient. The American Statistician, 26(2):31–33, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683460.

Counts:1979:HSS

[CS79] Richard Counts and Michael Sherrill. HYPERSAMS — a sample size computer program for the hypergeometric. The American Statistician, 33(2):91, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683232.

Canfield:1977:LE

[CTL^{+}77] Ronald V. Canfield, David L. Turner, Eugene Lehman, I. J. Good, Gottfried E. Noether, Benjamin J. Tepping, Bernard Most, Oliver D. Anderson, David A. Swanson, Roy H. Wampler, David Rothman, Russell V. Lenth, and Stephen Portnoy. Letters to the Editor. The American Statistician, 31(1):50–54, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682982.

Cutter:1976:SET

[Cut76] Gary R. Cutter. Some examples for teaching regression toward the mean from a sampling viewpoint. The American Statistician, 30(4):194–197, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683763.

DAgostino:1970:LEN

[D’A70] Ralph B. D’Agostino. Linear estimation of the normal distribution standard deviation. The American Statistician, 24(3):14–15, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682419.

DAgostino:1972:RBC

[D’A72] Ralph B. D’Agostino. Relation between the chi-squared and ANOVA tests for testing the equality of k independent dichotomous populations. The American Statistician, 26(3):30–32, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682863.

Darlington:1970:KRP

[Dar70] Richard B. Darlington. Is kurtosis really “peakedness?”. The American Statistician, 24(2):19–22, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681925.

David:1976:SIP

[Dav76] Martin David. Strategies for improving policy-relevant data on the poor. The American Statistician, 30(3):129–134, August 1976. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683859.

Dayton:1972:MCD

[Day72] C. Mitchell Dayton. A method for constructing data which illustrate a suppressor variable. The American Statistician, 26(5):36, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683784.

Dudewicz:1972:ELD

[DD72] Edward J. Dudewicz and Ronald E. Dann. Equally likely dice sums do not exist. The American Statistician, 26(4):41–42, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683210.

Dudewicz:1975:LE

[DDB^{+}75] Edward J. Dudewicz, Meyer Dwass, James P. Barrett, R. C. Hanumara, James R. Thompson, Clifford J. Maloney, Klaus Hinkelmann, Michael H. Kutner, and Norman L. Johnson. Letters to the Editor. The American Statistician, 29(2):108–110, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683281.

Deaton:1970:EML

[Dea70] Leonard W. Deaton. An example on maximum likelihood estimates. The American Statistician, 24(1):49, February 1970. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682299.

Deming:1971:FFS

[Dem71] W. Edwards Deming. Frederick Franklin Stephan, 1903–1971. The American Statistician, 25(4):47–48, October 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682929.

Deming:1972:PCM

[Dem72] W. Edwards Deming. P. C. Mahalanobis (1893–1972). The American Statistician, 26(4):49–50, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683213.

Deming:1975:PBA

[Dem75] W. Edwards Deming. On probability as a basis for action. The American Statistician, 29(4):146–152, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683482.

Desalvo:1971:SEF

[Des71a] Joseph S. Desalvo. Standard error of forecast in multiple regression: Proof of a useful result. The American Statistician, 25(4):32–34, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682924.

Desu:1971:OCI

[Des71b] M. M. Desu. Optimal confidence intervals of fixed width. The American Statistician, 25(2):27–29, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682004.

Dallal:1977:DDH

[DF77] Gerard Dallal and Katherine Finseth. Double dual histograms. The American Statistician, 31(1):39–41, February 1977. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682979.

Dickey:1975:LE

[DGK^{+}75] James Dickey, N. T. Gridgeman, M. C. S. Kingsley, I. J. Good, James E. Carlson, Daniel Gianola, Michael H. Kutner, and Steve Selvin. Letters to the Editor. The American Statistician, 29(3):131– 134, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683443.

Downs:1978:PCE

[DGK78] Tom Downs, Dennis C. Gilliland, and Leo Katz. Probability in a contested election. The American Statistician, 32(4):122–125, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682937.

Dykstra:1972:EDC

[DH72] Richard L. Dykstra and John E. Hewett. Examples of decompositions of chi-squared variables. The American Statistician, 26(4):42– 43, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683211.

Diamond:1975:CBS

[Dia75] James J. Diamond. [A course in Bayesian statistics]: Discussion. The American Statistician, 29(2):98–99, May 1975. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683275.

Driscoll:1975:AC

[DM75] M. F. Driscoll and D. Morse. Admissibility and completeness. The American Statistician, 29(2):93, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683273.

Deming:1970:LEb

[DMK^{+}70] W. Edwards Deming, Donald V. Mathusz, Scott A. Krane, Keith Kussmaul, Charles DeWitt Roberts, I. J. Good, Haynes Phillips, Charles J. Kowalski, Kenneth Mullen, Samuel H. Plotkin, Jason Millman, Paul Cieslak, S. Chatterjee, Nathan Mantel, Alan B. Forsythe, Joseph M. Moser, Ayala Cohen, Evelio Fabbroni, and Victor Chew. Letters to the Editor. The American Statistician, 24(4):46–53, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681316.

DAgostino:1973:EKS

[DN73] Ralph B. D’Agostino and Gottfried E. Noether. On the evaluation of the Kolmogorov statistic. The American Statistician, 27(2):81–82, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683963.

Doane:1976:AFC

[Doa76] David P. Doane. Aesthetic frequency classifications. The American Statistician, 30(4):181–183, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683757.

Doss:1973:MTC

[Dos73] D. C. Doss. Moments in terms of cumulants and vice versa. The American Statistician, 27(5):239–240, December 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/3087413.

Driscoll:1979:CPB

[DR79] Michael F. Driscoll and David A. Reynolds. A computer program for biased regression. The American Statistician, 33(3):160, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683827.

Drolette:1974:VTL

[Dro74] Margaret E. Drolette. The vanishing 2 × 2 table: Linking the hypergeometric, binomial and Poisson. The American Statistician, 28(3):102–103, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683036.

Davis:1978:LE

[DSMG78] John H. Davis, Norman Spector, J. E. Mann, and David Gilat. Letters to the Editor. The American Statistician, 32(2):76–77, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683624.

Deming:1970:LEa

[DTB^{+}70] W. Edwards Deming, H. Tamura, F. W. Bennett, Jr., I. H. LaValle, Eugene Dayhoff, William Knight, I. J. Good, L. R. Shenton, C. E. Stephenson, John D. Williams, Sudhir N. Dalal, Min-Te Chao, and M. N. Das. Letters to the Editor. The American Statistician, 24(2):35–39, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681931.

Duffett:1978:HM

[Duf78] Walter E. Duffett. Herbert Marshall, 1888–1977. The American Statistician, 32(2):75, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683623.

DuMouchel:1979:TSC

[DuM79] William H. DuMouchel. [teaching statistical computing using computer packages]: Comment. The American Statistician, 33 (1):30–31, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683066. See [Thi79a].

Duncan:1976:CFU

[Dun76] Joseph W. Duncan. Confidentiality and the future of the U.S. Statistical System. The American Statistician, 30(2):54–59, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683793.

Duncan:1977:DFP

[Dun77] Joseph W. Duncan. Developing “A Framework for Planning U.S. Federal Statistics, 1978–1989”. The American Statistician, 31 (3):107–113, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682955.

Dwass:1979:GBD

[Dwa79] Meyer Dwass. A generalized binomial distribution. The American Statistician, 33(2):86–87, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683228.

Deming:1978:LE

[DWTG78] W. Edwards Deming, P. David Wilson, James A. Tonascia, and William C. Guenther. Letters to the Editor. The American Statistician, 32(4):145–146, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682948.

Easterling:1973:PEB

[Eas73] R. G. Easterling. Point estimation and Bayesian inference. The American Statistician, 27(5):241, December 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/3087416.

Eaton:1978:YCM

[Eat78] P. W. Eaton. Yarnold’s criterion and minimum sample size. The American Statistician, 32(3):102–103, August 1978. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683353.

Eckler:1972:SCF

[Eck72] A. Ross Eckler. Some comments on “Federal Statistics” (Report of the President’s Commission on Federal Statistics — Volume I). The American Statistician, 26(2):11–14, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683451.

Ederer:1979:SRD

[Ede79] Fred Ederer. The Statistician’s role in developing a protocol for a clinical trial. The American Statistician, 33(3):116–119, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683811.

Eberhardt:1977:CTT

[EF77] Keith R. Eberhardt and Michael A. Fligner. A comparison of two tests for equality of two proportions. The American Statistician, 31(4):151–155, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683536.

Eisenhart:1970:AIS

[Eis70] Churchill Eisenhart. Anniversaries in 1970 of interest to statisticians. The American Statistician, 24(1):25–28, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682292.

Eisenhart:1975:SLP

[Eis75] Churchill Eisenhart. A supplementary list of publications of S. S. Wilks. The American Statistician, 29(1):25–27, February 1975. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683675.

Eisenhart:1979:TSS

[Eis79] Churchill Eisenhart. On the transition from “Student’s” z to “Student’s” t. The American Statistician, 33(1):6–10, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683058.

Evans:1976:RFI

[EJO76] I. G. Evans, D. A. Jones, and R. J. Owen. Relative frequency interpretation of Bayesian predictive distributions. The American Statistician, 30(3):145, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683862.

Emerson:1979:ALS

[ES79] John D. Emerson and Gary A. Simon. Another look at the sign test when ties are present: The problem of confidence intervals. The American Statistician, 33(3):140–142, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683818.

Evans:1973:MIC

[Eva73] Lawrence S. Evans. A mechanical interpretation of the coefficient of rank correlation and other analogies. The American Statistician, 27(2):79–81, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683962.

Faulkner:1977:RSS

[Fau77] J. Earl Faulkner. Review of The Statistician — special issues on statistics teaching in higher education. The American Statistician, 31(3):125–127, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682963.

Faulkner:1978:RST

[Fau78] J. Earl Faulkner. Review of statistical transparency sets. The American Statistician, 32(3):106–107, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683355.

Fox:1973:LE

[FDH^{+}73] Martin Fox, Herbert A. David, James R. Harvey, L. R. Shenton, Sigmund Amster, Alan Oppenheim, Elliot M. Cramer, Sylvia Wargon, Robert G. Easterling, and Lloyd S. Nelson. Letters to the Editor. The American Statistician, 27(2):90–93, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683969.

Federer:1978:SRS

[Fed78] Walter T. Federer. Some remarks on statistical education. The American Statistician, 32(4):117–121, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682936.

Feinberg:1971:TTT

[Fei71] William E. Feinberg. Teaching the Type I and Type II errors: The judicial process. The American Statistician, 25(3):30–32, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683322.

Federer:1973:SSD

[FF73] W. T. Federer and A. J. Federer. A study of statistical design publications from 1968 through 1971. The American Statistician, 27(4):160–163, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684047.

Ferber:1971:LE

[FGF^{+}71] Robert Ferber, Ernest W. Grove, Joseph L. Fleiss, Joe Engleman, Irving H. Siegel, Carl B. Bates, Y. Horiba, David S. Salsburg, Arthur F. Johnson, Lawrence L. Kupper, Barry C. Arnold, and Jin sheng Huang. Letters to the Editor. The American Statistician, 25(4):49–53, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682931.

Francis:1975:CCE

[FHV75] Ivor Francis, Richard M. Heiberger, and Paul F. Velleman. Criteria and considerations in the evaluation of statistical program packages. The American Statistician, 29(1):52–56, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683685.

Fienberg:1979:GMS

[Fie79] Stephen E. Fienberg. Graphical methods in statistics. The American Statistician, 33(4):165–178, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683729.

Fisher:1972:BBB

[Fis72] Lloyd Fisher. The behavior of Bayesians in bunches. The American Statistician, 26(5):19–20, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683775.

Fisher:1973:AAC

[Fis73] Lloyd Fisher. An alternative approach to Cochran’s theorem. The American Statistician, 27(3):109, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683632.

Folks:1970:SPP

[Fol70] J. Leroy Folks. Some prior probabilities on the future of statistics. The American Statistician, 24(5):10–13, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682443.

Fox:1979:TAW

[Fox79] John Fox. TAB: An APL workspace for the log-linear analysis of contingency tables. The American Statistician, 33(3):159–160, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683826.

Frane:1977:SSH

[Fra77] James W. Frane. Some statistical highlights of the BMDP series. The American Statistician, 31(3):131–133, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682965.

Freund:1971:SOR

[Fre71] R. J. Freund. Some observations on regressions with grouped data. The American Statistician, 25(3):29–30, June 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683321.

Friedman:1972:TJC

[Fri72] Herbert Friedman. Trial by jury: Criteria for convictions, jury size and Type I and Type II errors. The American Statistician, 26(2):21– 23, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683454.

Friess:1977:CSA

[Fri77] Seymour L. Friess. Contributions of statistics to the analysis of environmental health problems caused by pollutants. The American Statistician, 31(1):2–7, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682969.

Fleiss:1971:NPC

[FT71] Joseph L. Fleiss and Judith M. Tanur. A note on the partial correlation coefficient. The American Statistician, 25(1):43–45, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682214.

Faulkenberry:1979:SME

[FT79] G. David Faulkenberry and Robert D. Tortora. A study of measurement error suitable for a classroom example. The American Statistician, 33(1):19–22, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683061.

Fuchs:1979:PBI

[Fuc79] Camil Fuchs. Possible biased inferences in tests for average partial association. The American Statistician, 33(3):120–126, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683812.

Fuentes:1972:LAS

[Fue72] Sonia Pressman Fuentes. The law against sex discrimination in employment and its relationship to statistics. The American Statistician, 26(2):16–21, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683453.

Ferber:1974:LE

[FvFC^{+}74] Robert Ferber, F. W. von Fleckenstein, Harold F. Clark, Joe Davis, Harold L. Pazer, Lloyd A. Swanson, C. P. Quesenberry, Alan Ross, and J. A. John. Letters to the Editor. The American Statistician, 28(3):110–112, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683041.

Fligner:1976:SAS

[FW76] Michael A. Fligner and Douglas A. Wolfe. Some applications of sample analogues to the probability integral transformation and a coverage property. The American Statistician, 30(2):78–85, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683799.

Gallant:1975:NR

[Gal75] A. R. Gallant. Nonlinear regression. The American Statistician, 29(2):73–81, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683268.

Garfield:1974:WWR

[Gar74] Frank R. Garfield. Winfield W. Riefler, 1897–1974. The American Statistician, 28(4):159, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683344.

Gastwirth:1975:SME

[Gas75] Joseph L. Gastwirth. Statistical measures of earnings differentials. The American Statistician, 29(1):32–35, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683677.

Gastwirth:1977:PMP

[Gas77] Joseph L. Gastwirth. A probability model of a pyramid scheme. The American Statistician, 31(2):79–82, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683046.

Green:1974:DVS

[GD74] Richard D. Green and John P. Doll. Dummy variables and seasonality — a curio. The American Statistician, 28(2):60–62, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683603.

Gibbons:1976:LE

[GDS^{+}76] Jean D. Gibbons, Warren Dent, G. K. Shukla, G. A. Barnard, D. A. Sprott, Jean E. Weber, and David Monarchi. Letters to the Editor. The American Statistician, 30(3):150–152, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683865.

Geller:1978:SEW

[Gel78] Nancy L. Geller. Some examples of the weak and strong laws of large numbers for averages of mutually independent random variables. The American Statistician, 32(1):34–36, February 1978. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683474. See corrigenda

[Ano78e].

Gentleman:1977:APU

[Gen77] Jane F. Gentleman. It’s all a plot (using interactive computer graphics in teaching statistics). The American Statistician, 31 (4):166–175, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683541.

Gerber:1979:CPP

[Ger79] Hans U. Gerber. A characteristic property of the Poisson distribution. The American Statistician, 33(2):85–86, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683227.

Good:1977:LE

[GFM^{+}77] I. J. Good, Stephen E. Fienberg, James M. Malone, II, Maurice C. Bryson, Hugo C. Hamaker, R. M. Kozelka, Herbert F. Spirer, D. V. Lindley, and L. D. Phillips. Letters to the Editor. The American Statistician, 31(4):180–183, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683545.

Gibbons:1973:QE

[Gib73] Jean D. Gibbons. A question of ethics. The American Statistician, 27(2):72–76, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683960.

Giffler:1970:ISS

[Gif70] Milton Giffler. Industrial statistics of the State Planning Committee of the U.S.S.R. (Gosplan). The American Statistician, 24(2):10–16, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2681923.

Gilat:1977:MPF

[Gil77] David Gilat. Monotonicity of a power function: An elementary probabilistic proof. The American Statistician, 31(2):91–93, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683050.

Ginsberg:1971:IVS

[Gin71] William Ginsberg. Instrumental variables: Some generalizations. The American Statistician, 25(2):25–26, April 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682003.

Gapinski:1973:EAR

[GK73a] James H. Gapinski and T. Krishna Kumar. Errors in the application of recent multicollinearity tests. The American Statistician, 27(5):244, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087420.

Govindarajulu:1973:NAD

[GK73b] Z. Govindarajulu and J. H. Klotz. A note on the asymptotic distribution of the one-sample Kolmogorov–Smirnov statistic. The American Statistician, 27(4):164–165, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684048.

Gunderson:1979:BRB

[GK79] Norman G. Gunderson and Thomas R. Knapp. Review of the game Take a Chance. The American Statistician, 33(4):212–213, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683738.

Glick:1970:HPC

[Gli70] Ned Glick. Hijacking planes to Cuba: An up-dated version of the Birthday Problem. The American Statistician, 24(1):41–44, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682296.

Garside:1976:ATE

[GM76] G. R. Garside and C. Mack. Actual Type 1 error probabilities for various tests in the homogeneity case of the 2×2 contingency table. The American Statistician, 30(1):18–21, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682881.

Ghosh:1977:NAC

[GM77a] Malay Ghosh and Glen Meeden. On the non-attainability of Chebyshev bounds. The American Statistician, 31(1):35–36, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682977. See corri-

genda [Ano79g].

Groeneveld:1977:MMM

[GM77b] Richard A. Groeneveld and Glen Meeden. The mode, median, and mean inequality. The American Statistician, 31(3):120–121, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682959.

Gideon:1978:CTS

[GM78] Rudy A. Gideon and Dale E. Mueller. Computation of the twosample Smirnov statistics. The American Statistician, 32(4):136– 137, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2682941.

Godambe:1970:FSS

[God70] V. P. Godambe. Foundations of survey-sampling. The American Statistician, 24(1):33–38, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682294.

Godambe:1977:AB

[God77] V. P. Godambe. Allan Birnbaum, 1923–1976. The American Statistician, 31(4):178–179, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683544.

Gokhale:1972:FSS

[Gok72] D. V. Gokhale. A formula for sample size required for specified tolerance. The American Statistician, 26(5):21, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683776.

Goldstein:1973:AOE

[Gol73] Matthew Goldstein. An application of ordered expectations. The American Statistician, 27(1):30, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682905.

Goldfield:1975:AAC

[Gol75] Edwin D. Goldfield. ASA Advisory Committees to Federal Statistical Agencies: a report of the Committee on Committees of the American Statistical Association. The American Statistician, 29(2):103–105, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683278.

Good:1972:STP

[Goo72] I. J. Good. Statistics and Today’s problems. The American Statistician, 26(3):11–19, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682859.

Good:1973:WDF

[Goo73] I. J. Good. What are degrees of freedom? The American Statistician, 27(5):227–228, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/3087407.

Goodnight:1979:TSO

[Goo79] James H. Goodnight. A tutorial on the SWEEP operator. The American Statistician, 33(3):149–158, August 1979. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683825.

Gibbons:1978:BCE

[GOS78] Jean D. Gibbons, Ingram Olkin, and Milton Sobel. Baseball competitions — are enough games played? The American Statistician, 32(3):89–95, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683350.

Gibbons:1979:IRS

[GOS79] Jean D. Gibbons, Ingram Olkin, and Milton Sobel. An introduction to ranking and selection. The American Statistician, 33(4):185–195, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683731.

Gibbons:1975:VIM

[GP75] Jean D. Gibbons and John W. Pratt. P -values: Interpretation and methodology. The American Statistician, 29(1):20–25, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683674.

Gibbons:1974:LE

[GRB^{+}74] Jean D. Gibbons, David Rothman, Alvah C. Bittner, Jr., Paul I. Feder, Peter Bryant, Patrick E. O’Neil, Roy L. Streit, Harlley E. McKean, Frederick L. Newman, Dennis L. Young, P. S. Levy, S. H. Rosenberg, R. R. Hocking, and Alain Verghote. Letters to the Editor. The American Statistician, 28(1):35–40, February 1974. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683530.

Greene:1971:JSS

[Gre71] Jerome D. Greene. J. Stevens Stock, 1910–1971. The American Statistician, 25(4):48, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682930.

Green:1977:PIP

[Gre77] Bert F. Green. A practical interactive program for randomization tests of location. The American Statistician, 31(1):37–39, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682978.

Griffin:1979:BRB

[Gri79] John I. Griffin. Review of filmstrips–audiotape set: Introduction to Probability and Statistics. The American Statistician, 33(2):88–89, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683229.

Groeneveld:1971:NSS

[Gro71] Richard A. Groeneveld. A note on the sequential sign test. The American Statistician, 25(2):15–16, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681999.

Gurland:1971:SAU

[GT71] John Gurland and Ram C. Tripathi. A simple approximation for unbiased estimation of the standard deviation. The American Statistician, 25(4):30–32, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682923.

Gelfand:1974:CCA

[GT74] Alan E. Gelfand and D. L. Thomas. Confidence coefficients for approximate binomial interval estimates under hypergeometric models. The American Statistician, 28(2):52–54, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683599.

Guenther:1970:DSS

[Gue70] William C. Guenther. Determination of sample size for distributionfree tolerance limits. The American Statistician, 24(1):44–46, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682297.

Guenther:1971:NMV

[Gue71a] William C. Guenther. A note on the minimum variance unbiased estimate of the fraction of a normal distribution below a specification limit. The American Statistician, 25(2):18–20, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682001.

Guenther:1971:UCI

[Gue71b] William C. Guenther. Unbiased confidence intervals. The American Statistician, 25(1):51–53, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682218.

Guenther:1972:UIG

[Gue72] William C. Guenther. On the use of the incomplete gamma table to obtain unbiased tests and unbiased confidence intervals for the variance of a normal distribution. The American Statistician, 26(1):31–34, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682844.

Guenther:1975:SSF

[Gue75] William C. Guenther. A sample size formula for a non-central t test. The American Statistician, 29(3):120–121, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683437.

Guenther:1977:CFT

[Gue77a] William C. Guenther. Calculation of factors for tests and confidence intervals concerning the ratio of two normal variances. The American Statistician, 31(4):175–177, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683542.

Guenther:1977:DCP

[Gue77b] William C. Guenther. Desk calculation of probabilities for the distribution of the sample correlation coefficient. The American Statistician, 31(1):45–48, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682981.

Guenther:1977:PSSb

[Gue77c] William C. Guenther. Power and sample size determination when the alternative hypotheses are given in quantiles. The American Statistician, 31(3):117–118, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682957.

Guenther:1977:PSSa

[Gue77d] William C. Guenther. Power and sample size for approximate chisquare tests. The American Statistician, 31(2):83–85, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683047.

Guenther:1977:SPA

[Gue77e] William C. Guenther. Some probabilities available from desk calculators and their relation to tables. The American Statistician, 31(1):41–45, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682980.

Guenther:1978:SEF

[Gue78a] William C. Guenther. Some easily found minimum variance unbiased estimators. The American Statistician, 32(1):29–34, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683473.

Guenther:1978:SRR

[Gue78b] William C. Guenther. Some remarks on the runs test and the use of the hypergeometric distribution. The American Statistician, 32(2): 71–73, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683620.

Guenther:1979:UNA

[Gue79] William C. Guenther. The use of noncentral F approximations for calculation of power and sample size. The American Statistician, 33(4):209–210, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683736.

Gujarati:1970:UDVb

[Guj70a] Damodar Gujarati. Use of dummy variables in testing for equality between sets of coefficients in linear regressions: a generalization.

The American Statistician, 24(5):18–22, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682446.

Gujarati:1970:UDVa

[Guj70b] Damodar Gujarati. Use of dummy variables in testing for equality between sets of coefficients in two linear regressions: a note. The American Statistician, 24(1):50–52, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682300.

Hoaglin:1975:RCB

[HA75] David C. Hoaglin and David F. Andrews. The reporting of computation-based results in statistics. The American Statistician, 29(3):122–126, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683438.

Hader:1973:IMR

[Had73] R. J. Hader. An improper method of randomization in experimental design. The American Statistician, 27(2):82–84, April 1973. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683964.

Han:1971:DRV

[Han71] Chien-Pai Han. Dependence of random variables. The American Statistician, 25(4):35, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682925.

Hanushek:1974:EER

[Han74] Eric A. Hanushek. Efficient estimators for regressing regression coefficients. The American Statistician, 28(2):66–67, May 1974. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683605.

Han:1979:AME

[Han79] Chien Pai Han. Alternative methods of estimating the likelihood ratio in classification of multivariate normal observations. The American Statistician, 33(4):204–206, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683733.

Harter:1970:MCP

[Har70] H. Leon Harter. Multiple comparison procedures for interactions. The American Statistician, 24(5):30–32, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682450.

Harris:1972:NDP

[Har72a] Lawrence Harris. A note on the decision process of the official statistician. The American Statistician, 26(2):23–24, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683455.

Harvey:1972:EMT

[Har72b] James R. Harvey. On expressing moments in terms of cumulants and vice versa. The American Statistician, 26(4):38–39, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683208. See cor-

rections [Ano73n].

Harshbarger:1974:EHD

[Har74] Boyd Harshbarger. An example of how a designed experiment saved research workers from false recommendations. The American Statistician, 28(4):128–129, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683336.

Harville:1975:ERW

[Har75] David A. Harville. Experimental randomization: Who needs it? The American Statistician, 29(1):27–31, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683676.

Harvey:1976:APG

[Har76] A. C. Harvey. An alternative proof and generalization of a test for structural change. The American Statistician, 30(3):122–123, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683857.

Hauser:1973:SP

[Hau73] Philip M. Hauser. Statistics and politics. The American Statistician, 27(2):68–71, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683959.

Hawkins:1974:NDT

[Haw74] Douglas M. Hawkins. A note on the distribution of times reported by a low resolution interval timer. The American Statistician, 28(2): 56–57, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683600.

Hawkins:1975:NNI

[Haw75] Douglas M. Hawkins. From the noncentral t to the normal integral. The American Statistician, 29(1):42–43, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683681.

Heiberger:1979:TSC

[Hei79] Richard M. Heiberger. [teaching statistical computing using computer packages]: Comment. The American Statistician, 33(1):32– 33, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/

2683067. See [Thi79a].

Herbach:1972:PPD

[Her72] Leon H. Herbach. Properties of the Pascal distribution without recourse to the binomial theorem with negative exponent. The American Statistician, 26(4):39–41, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683209. See corrections [Ano73o].

Hoyt:1970:LE

[HGS^{+}70] John P. Hoyt, Ernest W. Grove, H. Fairfield Smith, M. A. Hamdan, A. L. Wilson, Seymour Geisser, H. Weiler, Shui-Mu Huang, B. M. Bennett, Richard M. Brugger, and Edward E. Cureton. Letters to the Editor. The American Statistician, 24(1):58–61, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682303.

Holloway:1976:LE

[HHN^{+}76] Clark Holloway, David R. Howes, G. Nagase, K. S. Banerjee, Nathan Mantel, Lawrence L. Kupper, Anthony J. McMichael, Robert Spirtas, J. H. Maindonald, John J. Gart, M. R. Heyworth, W. K. Chiu, K. Lam, Lawrence Klimko, Paul Nelson, R. R. Hocking, and F. M. Speed. Letters to the Editor. The American Statistician, 30(4):200–206, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683766.

Higgins:1975:SSI

[Hig75] James J. Higgins. Some surface integral techniques in statistics. The American Statistician, 29(1):43–46, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683682.

Hildebrand:1971:KMB

[Hil71] David K. Hildebrand. Kurtosis measures bimodality? The American Statistician, 25(1):42–43, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682213.

Hendrickson:1973:LE

[HJN^{+}73] John A. Hendrickson, Jr., Arthur F. Johnson, Gottfried E. Noether, Bernard Harris, Andrew P. Soms, Daniel L. Solomon, R. W. Farebrother, Richard J. Foote, J. Edward Ely, Clifford J. Maloney, Donald A. Berry, John Kinney, and George A. F. Seber. Letters to the Editor. The American Statistician, 27(4):191–196, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684055.

Hettmansperger:1974:GRN

[HM74] Thomas P. Hettmansperger and Joseph W. McKean. A graphical representation for non-parametric inference. The American Statistician, 28(3):100–102, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683035.

Huang:1978:CPU

[HMC78] Chung-Liang Huang, Leonard C. Moon, and Hui S. Chang. A computer program using the Box–Cox transformation technique for the specification of functional form. The American Statistician, 32(4):144, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682947.

Hocking:1973:DTW

[Hoc73] R. R. Hocking. A discussion of the two-way mixed model. The American Statistician, 27(4):148–152, October 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684043.

Hochberg:1974:DRG

[Hoc74] Yosef Hochberg. The distribution of the range in general balanced models. The American Statistician, 28(4):137–138, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683339.

Hogg:1972:SE

[Hog72] Robert V. Hogg. On statistical education. The American Statistician, 26(3):8–11, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682858.

Hogg:1979:SRO

[Hog79] Robert V. Hogg. Statistical robustness: one view of its use in applications today. The American Statistician, 33(3):108–115, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683810.

Holshouser:1974:CPV

[Hol74] Eugene C. Holshouser. Computing the posterior variance in the normal case. The American Statistician, 28(1):23–24, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683526.

Houghton:1973:CCA

[Hou73] E. L. Houghton. Course charts: An application. The American Statistician, 27(1):25–27, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682901.

Hoyt:1972:GFE

[Hoy72] John P. Hoyt. Generating functions in elementary probability theory. The American Statistician, 26(3):45–46, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682868.

Hamdan:1971:LE

[HPG^{+}71] M. A. Hamdan, John W. Pratt, Peter Gottlieb, I. J. Good, M. A. Hamdan, S. G. Carmer, W. M. Walker, T. J. Valentine, Ralph B. D’Agostino, A. M. Kshirsagar, I. Gwyn Evans, D. G. Kabe, Edward E. Cureton, J. R. Rutherford, J. K. Sharma, and James R. Harvey. Letters to the Editor. The American Statistician, 25 (5):56–63, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686089.

Hartley:1971:FSS

[HR71] H. O. Hartley and J. N. K. Rao. Foundations of survey sampling (a Don Quixote tragedy). The American Statistician, 25(1):21–27, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682207.

Heath:1976:FTE

[HS76a] David Heath and William Sudderth. De Finetti’s Theorem on exchangeable variables. The American Statistician, 30(4):188–189, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683760.

Hurlburt:1976:DRS

[HS76b] Russell T. Hurlburt and Douglas K. Spiegel. Dependence of F ratios sharing a common denominator mean square. The American Statistician, 30(2):74–78, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683798.

Hsiao:1972:DRC

[Hsi72] Frank S. T. Hsiao. The diagrammatical representation of confidence-interval estimation and hypothesis testing. The American Statistician, 26(5):28–29, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683779.

Huang:1976:BRC

[Hua76] Cliff J. Huang. Bias in routine calculation of the variance of OLS estimators when the disturbances are autocorrelated. The American Statistician, 30(2):70–72, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683796.

Hunter:1972:ICE

[Hun72] Jeffrey J. Hunter. Independence, conditional expectation, and zero covariance. The American Statistician, 26(5):22–24, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683777.

Hunter:1977:SIA

[Hun77] W. G. Hunter. Some ideas about teaching design of experiments, with 2^{5} examples of experiments conducted by students. The American Statistician, 31(1):12–17, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682971.

Hoaglin:1978:MRA

[HW78] David C. Hoaglin and Roy E. Welsch. The hat matrix in regression and ANOVA. The American Statistician, 32(1):17–22, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683469. See

corrigenda [Ano78f].

Inglis:1978:BRB

[IK78] James Inglis and Thomas R. Knapp. Book review: Statistics at a Glance. The American Statistician, 32(3):107, August 1978. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683356.

Iskow:1979:EMI

[IS79] Lawrence I. Iskow and Randall K. Spoeri. EASYMAP: a machineindependent line printer statistical mapping system. The American Statistician, 33(4):223–224, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683744.

Janardan:1971:NPC

[Jan71] K. G. Janardan. Note on the problem of class-size distributions. The American Statistician, 25(1):45–46, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682215.

Janardan:1973:AEH

[Jan73] K. G. Janardan. On an alternative expression for the hypergeometric moment generating function. The American Statistician, 27(5):242, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087418.

Jayachandran:1970:DDT

[JB70] Toke Jayachandran and D. R. Barr. On the distribution of a difference of two scaled chi-square random variables. The American Statistician, 24(5):29–30, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682449.

Johnson:1976:TVP

[JK76] Norman L. Johnson and Samuel Kotz. Two variants of Polya’s urn models (towards a rapprochement between combinatorics and probability theory). The American Statistician, 30(4):186–188, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683759.

Johnson:1973:CRR

[JKS73] Dallas E. Johnson, John D. Kubicek, and Stephen M. Stigler. Is “complete” really the right name? The American Statistician, 27(5):240–241, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087415.

Jensen:1976:PEH

[JM76] Bartell C. Jensen and James B. McDonald. A pedagogical example of heteroskedasticity and autocorrelation. The American Statistician, 30(4):192–193, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683762.

Johnson:1973:DSM

[Joh73] Dallas E. Johnson. A derivation of Scheffe’s S-method by maximizing a quadratic form. The American Statistician, 27(1):27–29, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682903.

Johnson:1974:EMT

[Joh74a] B. R. Johnson. An effective method for teaching the Lehmann– Scheffe technique for obtaining a minimal sufficient statistic. The American Statistician, 28(2):59–60, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683602.

Johnson:1974:PFE

[Joh74b] D. Gale Johnson. Population, food and economic adjustment. The American Statistician, 28(3):89–93, August 1974. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683031.

Joiner:1974:CEN

[Joi74] Brian L. Joiner. Critical evaluation of numerical data. The American Statistician, 28(1):14–16, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683521.

Joshi:1970:CCB

[Jos70] S. W. Joshi. Construction of certain bivariate distributions. The American Statistician, 24(2):32, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681929.

Kshirsagar:1975:NET

[KA75] A. M. Kshirsagar and E. Arseven. A note on the equivalency of two discrimination procedures. The American Statistician, 29(1):38–39, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683679.

Kabak:1970:WMA

[Kab70] Irwin W. Kabak. Wildlife management: An application of a finite Markov chain. The American Statistician, 24(5):27–29, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682448.

Kappenman:1979:UCC

[Kap79] Russell F. Kappenman. On the use of a certain conditional distribution to obtain unconditional distributional results. The American Statistician, 33(1):23–24, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683062.

Katz:1975:PCI

[Kat75a] Leo Katz. Presentation of a confidence interval estimate as evidence in a legal proceeding. The American Statistician, 29(4):138–142, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683480.

Katz:1975:DDC

[Kat75b] Leon Katz. A direct development of the correlation coefficient. The American Statistician, 29(4):170, November 1975. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683488.

Kendall:1973:LE

[KBE^{+}73] Barbie Kendall, Irwin D. J. Bross, Laszlo Engelman, Boyd Z. Palmer, E. Whyte, John Wiorkowski, Nathan Mantel, Marvin S. Margolis, J. Phillip Cooper, Y. S. Sathe, I. J. Good, George Zyskind, and Paul A. Johnson. Letters to the Editor. The American Statistician, 27(1):39–44, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682908.

Kopitzke:1975:LSP

[KBG75] R. Kopitzke, T. J. Boardman, and F. A. Graybill. Least squares programs — a look at the square root procedure. The American Statistician, 29(1):64–66, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683688.

Kish:1971:LE

[KBR^{+}71] Leslie Kish, Irwin D. J. Bross, Fred L. Ramsey, V. P. Godambe, John P. Hoyt, Edward J. Carney, Sam Kemel, Ram C. Dahiya, J. Stewart, and Hardeo Sahai. Letters to the Editor. The American Statistician, 25(2):36–41, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682007.

Knight:1975:LE

[KBT^{+}75] William Knight, James K. Brewer, Irene M. Trawinski, H. Paul Greenough, Francisco J. Samaniego, A. K. Gupta, David Rothman, D. W. Turner, Douglas M. Hawkins, Marvin J. Karson, Taka Ashikaga, and L. R. Kenward. Letters to the Editor. The American Statistician, 29(4):171–175, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683489.

Keats:1974:SPG

[Kea74] John B. Keats. Some Pascal generalizations. The American Statistician, 28(1):24–25, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683527.

Kempthorne:1975:GZ

[Kem75] Oscar Kempthorne. George Zyskind, 1929–1974. The American Statistician, 29(2):106–107, May 1975. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683280.

Kendall:1976:SPC

[Ken76] Maurice Kendall. Statisticians — production and consumption. The American Statistician, 30(2):49–53, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683792.

Keyfitz:1970:FPO

[Key70] Nathan Keyfitz. Finding probabilities from observed rates or how to make a life table. The American Statistician, 24(1):28–33, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682293.

Keyfitz:1979:IATa

[Key79a] Nathan Keyfitz. Information and allocation: Two uses of the 1980 census. The American Statistician, 33(2):45–50, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683218. See comments [Nis79,

Rob79] and rejoinder [Key79b].

Keyfitz:1979:IATb

[Key79b] Nathan Keyfitz. [information and allocation: Two uses of the 1980 census]: Rejoinder. The American Statistician, 33(2):55–56, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683221. See

[Key79a].

Kotz:1973:SDS

[KJ73] Samuel Kotz and Norman L. Johnson. Statistical distributions: a survey of the literature, trends and prospects. The American Statistician, 27(1):15–17, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682898.

Ku:1974:LMC

[KK74] Harry H. Ku and Solomon Kullback. Loglinear models in contingency table analysis. The American Statistician, 28(4):115–122, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683334.

Kleinbaum:1978:BRB

[KK78] David G. Kleinbaum and Roy R. Kuebler. Review of two films on statistics. The American Statistician, 32(2):73–74, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683621.

Knight:1973:CMP

[Kni73] William Knight. The compartmental model paradox. The American Statistician, 27(5):246, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/3087425.

Kotlarski:1975:SDU

[Kot75] Ignacy I. Kotlarski. Some deficiencies of using moment generating functions. The American Statistician, 29(3):127–128, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683439.

Kowalski:1973:NNB

[Kow73] Charles J. Kowalski. Non-normal bivariate distributions with normal marginals. The American Statistician, 27(3):103–106, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683630.

Kronmal:1979:AMG

[KP79] Richard A. Kronmal and Arthur V. Peterson, Jr. On the alias method for generating random variables from a discrete distribution. The American Statistician, 33(4):214–218, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683739.

Kruskal:1971:MSS

[Kru71] William H. Kruskal. Mathematical sciences and social sciences: Excerpts from the report of a Panel of the Behavioral and Social Sciences Survey. The American Statistician, 25(1):27–31, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682208.

Krutchkoff:1972:EBE

[Kru72] Richard G. Krutchkoff. Empirical Bayes estimation. The American Statistician, 26(5):14–16, December 1972. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683773.

Kruskal:1974:US

[Kru74] William Kruskal. The ubiquity of statistics. The American Statistician, 28(1):3–6, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683518.

Kaiser:1972:PDT

[KS72] Henry F. Kaiser and Wilhelmine Stefansky. A Polya distribution for teaching. The American Statistician, 26(3):40–43, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682866.

Kshirsagar:1971:BDM

[Ksh71] A. M. Kshirsagar. Bias due to missing plots. The American Statistician, 25(1):47–50, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682217.

Koop:1974:LE

[KsLG^{+}74] J. C. Koop, Jack Chao sheng Lee, Fernando Garagorry, Jack Mardekian, Farhad Mehran, James C. Impara, Irwin D. J. Bross, John P. Hoyt, Richard L. Haden, Don Self, Marvin A. Johnston, A. R. Sampson, David Hinkley, and I. J. Good. Letters to the Editor. The American Statistician, 28(2):73–77, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683609.

Koehn:1975:SIS

[KT75] Uwe Koehn and D. L. Thomas. On statistics independent of a sufficient statistic: Basu’s lemma. The American Statistician, 29 (1):40–42, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683680.

Kutner:1974:HTL

[Kut74] Michael H. Kutner. Hypothesis testing in linear models (Eisenhart Model I). The American Statistician, 28(3):98–100, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683034.

Kymn:1970:STP

[Kym70] Kern O. Kymn. Statistical test of a partial regression coefficient under zero population correlation. The American Statistician, 24(2):30–31, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681928.

Lawrance:1976:CPC

[Law76] A. J. Lawrance. On conditional and partial correlation. The American Statistician, 30(3):146–149, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683864.

Lawrance:1979:PMC

[Law79] A. J. Lawrance. Partial and multiple correlation for time series. The American Statistician, 33(3):127–130, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683813.

Lachenbruch:1971:SDP

[LB71] Peter A. Lachenbruch and Donna R. Brogan. Some distributions on the positive real line which have no moments. The American Statistician, 25(1):46–47, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682216.

Larson:1972:NPR

[LB72] H. J. Larson and D. R. Barr. A note on pooling regression estimates. The American Statistician, 26(5):35, December 1972. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683782.

Lee:1978:ETS

[Lee78] S. Keith Lee. An example for teaching some basic concepts in multidimensional contingency table analysis. The American Statistician, 32(2):69–71, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683619.

Lehman:1971:RST

[Leh71] John W. Lehman. 1970 report of the Secretary–Treasurer. The American Statistician, 25(3):58–60, June 1971. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683330.

Lehman:1972:RST

[Leh72] John W. Lehman. 1971 report of the Secretary–Treasurer. The American Statistician, 26(3):66–71, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682872.

Lehman:1973:RST

[Leh73] John W. Lehman. 1972 report of the Secretary–Treasurer. The American Statistician, 27(3):129–133, June 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683642.

Lentner:1973:SPF

[Len73] Marvin Lentner. On the sum of powers of the first N integers. The American Statistician, 27(2):87, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683967.

Lessing:1973:AEH

[Les73] Ronald Lessing. An alternative expression for the hypergeometric moment generating function. The American Statistician, 27(3):115, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683636.

Levene:1974:HH

[Lev74] Howard Levene. Harold Hotelling, 1895–1973. The American Statistician, 28(2):71–73, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683608.

Levy:1979:PCA

[Lev79] Kenneth J. Levy. Pairwise comparisons associated with the K independent sample median test. The American Statistician, 33(3):138– 139, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683817.

Leaverton:1972:LE

[LFB^{+}72] Paul E. Leaverton, R. J. Freund, D. E. Barton, Bruce R. Boals, Lee R. Abramson, James L. Bicksler, T. Cacoullos, I. J. Good, Masoodul Haq, James R. Harvey, Joel Huber, D. M. Olsson, Tony Lancaster, S. R. Searle, R. Choudary Hanumara, Edward F. Gocka, James R. Maar, and Richard M. Brugger. Letters to the Editor. The American Statistician, 26(2):42–52, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683463.

Lurie:1972:MGO

[LH72] D. Lurie and H. O. Hartley. Machine-generation of order statistics for Monte Carlo computations. The American Statistician, 26(1):26–27, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682842.

Light:1975:CBS

[Lig75] Richard J. Light. [A course in Bayesian statistics]: Discussion. The American Statistician, 29(2):99–101, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683276.

Leung:1975:NGR

[LL75] Chi-Kun Leung and Kin Lam. A note on the geometric representation of the correlation coefficients. The American Statistician, 29(3):128–130, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683440.

Li:1975:PCT

[LMR75] C. C. Li, Sati Mazumdar, and B. Raja Rao. Partial correlation in terms of path coefficients. The American Statistician, 29(2):89–90, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683271.

Lord:1975:AHT

[Lor75] Frederic M. Lord. Automated hypothesis tests and standard errors for nonstandard problems. The American Statistician, 29(1):56–59, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683686.

Lindley:1976:IBP

[LP76] D. V. Lindley and L. D. Phillips. Inference for a Bernoulli process (a Bayesian view). The American Statistician, 30(3):112–119, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683855.

Lucas:1973:LDB

[Luc73] James M. Lucas. Large differences between partial correlation coefficients. The American Statistician, 27(2):77–78, April 1973. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683961.

Leinhardt:1979:TRE

[LW79] Samuel Leinhardt and Stanley S. Wasserman. Teaching regression: An exploratory approach. The American Statistician, 33 (4):196–203, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683732.

Makridakis:1979:CIA

[Mak79] Spyros Makridakis. CLT: An interactive approach for illustrating the Central Limit Theorem. The American Statistician, 33(2):90, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683230.

Malkiel:1978:PFE

[Mal78] Burton G. Malkiel. Problems with the federal economic statistical system and some alternatives for improvement. The American Statistician, 32(3):81–88, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683349.

Mallows:1979:RMS

[Mal79] Colin L. Mallows. Robust methods — some examples of their use. The American Statistician, 33(4):179–184, November 1979. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683730.

Mantel:1972:PCD

[Man72] Nathan Mantel. A property of certain distributions. The American Statistician, 26(2):29–30, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683459.

Mantel:1973:CFE

[Man73a] Nathan Mantel. A characteristic function exercise. The American Statistician, 27(1):31, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682906.

Mantell:1973:ELR

[Man73b] Edmund H. Mantell. Exact linear restrictions on parameters in the classical linear regression model. The American Statistician, 27(2):86–87, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683966.

Mandel:1974:SST

[Man74] B. J. Mandel. Statistical science training for government employees. The American Statistician, 28(1):7–9, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683519.

Mangold:1975:NFV

[Man75] Donald C. Mangold. A new formula for the variance of a systematic sample mean using serial correlations. The American Statistician, 29(3):114–115, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683434.

Mantel:1976:TD

[Man76a] Nathan Mantel. Tails of distributions. The American Statistician, 30(1):14–17, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682880.

Mantel:1976:TIT

[Man76b] Nathan Mantel. Transformations and the income tax. The American Statistician, 30(3):124–129, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683858.

Marshall:1970:SDM

[Mar70] Clifford W. Marshall. A simple derivation of the mean and variance of the number of runs in an ordered sample. The American Statistician, 24(4):27–28, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681311.

Martin:1974:WCN

[Mar74] Margaret E. Martin. The work of the Committee on National Statistics. The American Statistician, 28(3):104–107, August 1974. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683038.

Martin:1976:RAC

[Mar76] Margaret E. Martin. Report on activities: Committee on National Statistics. The American Statistician, 30(1):21–23, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682882.

Margolis:1979:PPE

[Mar79a] Marvin S. Margolis. Perpendicular projections and elementary statistics. The American Statistician, 33(3):131–135, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683814.

Marquardt:1979:SCI

[Mar79b] Donald W. Marquardt. Statistical consulting in industry. The American Statistician, 33(3):102–107, August 1979. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683809.

Matis:1973:PCM

[Mat73] James H. Matis. A paradox on compartmental models with Poisson immigration. The American Statistician, 27(2):84–85, April 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683965.

Maxwell:1974:EVO

[Max74] E. A. Maxwell. Estimating variances from one or two measurements on each sample. The American Statistician, 28(3):96–97, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683033.

Manson:1975:OZM

[MBG75] A. R. Manson, A. J. Barr, and J. H. Goodnight. Optimum zeromemory strategy and exact probabilities for 4-deck blackjack. The American Statistician, 29(2):84–88, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683270. See corrections [Ano75e].

McCafferty:1975:RHB

[McC75] James A. McCafferty. Ronald H. Beattie, 1903–1974. The American Statistician, 29(3):131, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683442.

McVay:1978:HLL

[McV78] F. E. McVay. Henry Laurence Lucas, Jr., 1916–1977. The American Statistician, 32(1):38, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683476.

Morgan:1972:LE

[MDC^{+}72] Joseph A. Morgan, Angus S. Deaton, Elliot M. Cramer, John Bibby, John J. Wiorkowski, Roger H. Moore, J. P. O’Neill, and Hal R. Varian. Letters to the Editor. The American Statistician, 26(3):62– 65, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682871.

Malone:1971:LE

[MDH^{+}71] J. M. Malone, II, Robyn M. Dawes, Charles C. Holt, Purnell Benson, S. S. Lewis, Samuel H. Plotkin, W. Edwards Deming, H. Leon Harter, Richard F. Potthoff, Giovanni Girone, Hubert Tolman, Anil P. Gore, Nader Fergany, Wei Shih, Thomas J. Valentine, and Hardeo Sahai. Letters to the Editor. The American Statistician, 25(1):56–63, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682220.

Meeker:1978:TUO

[Mee78] William Q. Meeker, Jr. TSERIES — a user-oriented computer program for time series analysis. The American Statistician, 32(3):111– 112, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683361.

Mendelssohn:1970:PPS

[Men70] Rudolph C. Mendelssohn. The principles of processing statistical data. The American Statistician, 24(4):15–19, October 1970. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681308.

Menezes:1971:DVM

[Men71] O. J. Menezes. The dummy variable method for jointly estimating seasonal indexes and regression on non-seasonal variables. The American Statistician, 25(5):32–36, December 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686081.

Mendenhall:1979:RPW

[Men79] William Mendenhall. Review of the Proceedings of the Workshop on the Teaching of Statistics. The American Statistician, 33(1):26, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683064.

Merriam:1971:AMV

[Mer71] Ida C. Merriam. ASA membership views on annual meetings. The American Statistician, 25(5):13–17, December 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686075.

Meyer:1974:CL

[Mey74] R. M. Meyer. On correlation and linearity. The American Statistician, 28(3):103, August 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683037.

Minton:1977:OCS

[MF77] P. D. Minton and R. J. Freund. Organization for the conduct of statistical activities in colleges and universities. The American Statistician, 31(3):113–117, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682956.

Mantel:1971:PAC

[MH71] Nathan Mantel and Benjamin F. Hankey. Programmed analysis of a 2 × 2 contingency table. The American Statistician, 25(5):40–44, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2686083.

Mantel:1975:ORC

[MH75] Nathan Mantel and Benjamin F. Hankey. The odds ratios of a 2 × 2 contingency table. The American Statistician, 29(4):143–145, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683481.

Maghsoodloo:1976:RSM

[MH76] S. Maghsoodloo and J. N. Hool. On response surface methodology and its computer-assisted teaching. The American Statistician, 30(3):140–144, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683861.

Meeker:1975:CPE

[MHF75] William Q. Meeker, Jr., Gerald J. Hahn, and Paul I. Feder. A computer program for evaluating and comparing experimental designs and some applications. The American Statistician, 29(1):60–64, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683687.

Meeker:1977:NBE

[MHF77] William Q. Meeker, Jr., Gerald J. Hahn, and Paul I. Feder. New bias evaluation features of EXPLOR — a program for assessing experimental design properties. The American Statistician, 31(2): 95–96, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683053.

Moussa-Hamouda:1977:ETN

[MHL77] Effat Moussa-Hamouda and Fred C. Leone. The effect of trimming and normality assumptions on the efficiency of the O-BLUE estimators. The American Statistician, 31(4):158–160, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683538.

Makridakis:1974:IFS

[MHW74] Spyros Makridakis, Anne Hodgsdon, and Steven C. Wheelwright. An interactive forecasting system. The American Statistician, 28 (4):153–158, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683343.

Miller:1978:IFT

[Mil78] John J. Miller. The inverse of the Freeman–Tukey double arcsine transformation. The American Statistician, 32(4):138, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682942.

Moran:1971:ISS

[MN71] P. A. P. Moran and J. Neyman. The IASPS and its satellite symposia. The American Statistician, 25(1):15–18, February 1971. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682205.

Monlezun:1979:TDP

[Mon79] Charles J. Monlezun. Two-dimensional plots for interpreting interactions in the three-factor analysis of variance model. The American Statistician, 33(2):63–69, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683223.

Morrison:1973:SRW

[Mor73] Donald G. Morrison. Some results for waiting times with an application to survey data. The American Statistician, 27(5):226–227, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/3087406.

Morris:1975:CIO

[Mor75] Wilson D. Morris, Jr. Combinations of indistinct objects. The American Statistician, 29(1):50–51, February 1975. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683684.

Morrison:1978:PMF

[Mor78] Donald G. Morrison. A probability model for forced binary choices. The American Statistician, 32(1):23–25, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683470.

Mosteller:1970:PRJ

[Mos70] Frederick Mosteller. Progress report of the Joint Committee of the American Statistical Association and the National Council of Teachers of Mathematics. The American Statistician, 24(3):8–12, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682417.

Moss:1972:RB

[Mos72] Milton Moss. Raymond T. Bowman, 1902–1971. The American Statistician, 26(1):45–46, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682848.

Moser:1976:RCS

[Mos76] Claus Moser. The role of the Central Statistical Office in assisting public policy makers. The American Statistician, 30(2):59–67, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683794.

Moses:1979:EST

[Mos79a] Lincoln E. Moses. Early steps toward a National Energy Information System. The American Statistician, 33(3):97–101, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683808.

Moss:1979:ER

[Mos79b] Milton Moss. Ernest Rubin, 1915–1978. The American Statistician, 33(4):208, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683735.

McBride:1979:LE

[MRT^{+}79] Noel A. McBride, Markley Roberts, Lazare Teper, Burton G. Malkiel, John P. Buonaccorsi, Gary O. Zerbe, Sanford Weisberg, James R. Beniger, Dorothy L. Robyn, and Claude Angers. Letters to the Editor. The American Statistician, 33(3):161–164, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683828.

Marquardt:1975:RRP

[MS75] Donald W. Marquardt and Ronald D. Snee. Ridge regression in practice. The American Statistician, 29(1):3–20, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683673.

Michael:1976:GRV

[MSH76] John R. Michael, William R. Schucany, and Roy W. Haas. Generating random variates using transformations with multiple roots. The American Statistician, 30(2):88–90, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683801.

McGill:1978:VBP

[MTL78] Robert McGill, John W. Tukey, and Wayne A. Larsen. Variations of Box plots. The American Statistician, 32(1):12–16, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683468.

Mullet:1972:GIS

[Mul72] Gary M. Mullet. A graphical illustration of simple (total) and partial regression. The American Statistician, 26(5):25–27, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683778.

Mullet:1977:SPN

[Mul77] Gary M. Mullet. Simeon Poisson and the National Hockey League. The American Statistician, 31(1):8–12, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682970.

Munford:1977:NUA

[Mun77] A. G. Munford. A note on the uniformity assumption in the Birthday Problem. The American Statistician, 31(3):119, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682958.

Makridakis:1978:NSR

[MW78] Spyros Makridakis and Steven C. Wheelwright. The new SIBYL/ RUNNER Interactive Forecasting Package (1978). The American Statistician, 32(3):109–110, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683359.

Namias:1978:RTE

[Nam78] Jean Namias. Review of three exercises in data presentation. The American Statistician, 32(2):74, May 1978. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683622.

Narula:1974:CLE

[Nar74] Subhash C. Narula. Correction: Letters to the Editor. The American Statistician, 28(4):166, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683346.

Nelson:1972:SMR

[Nel72] Wayne Nelson. Statistical methods for the ratio of two multinomial proportions. The American Statistician, 26(3):22–27, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682861.

Nelson:1976:IAM

[Nel76] Charles R. Nelson. The interpretation of R^{2} in autoregressivemoving average time series models. The American Statistician, 30(4):175–180, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683756.

Nelson:1977:UTN

[Nel77] Lloyd S. Nelson. Using tenth and ninetieth percentiles and the mode to characterize lognormal distributions. The American Statistician, 31(1):26–30, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682974.

Nelson:1979:BRB

[Nel79] Lloyd S. Nelson. Review of videotape course on manufacturing quality control. The American Statistician, 33(3):145–147, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683822.

Novick:1979:CAD

[NHC79] Melvin R. Novick, Robert M. Hamer, and James J. Chen. The Computer-Assisted Data Analysis (CADA) Monitor (1978). The American Statistician, 33(4):219–220, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683740.

Nisselson:1979:IAT

[Nis79] Harold Nisselson. [information and allocation: Two uses of the 1980 census]: Comment. The American Statistician, 33(2):50–52, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683219. See

[Key79a].

Neudecker:1973:WAL

[NJ73] H. Neudecker and Dallas E. Johnson. Who’s afraid of Lagrange multipliers? The American Statistician, 27(5):243–244, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/3087419.

Noether:1972:DFC

[Noe72] Gottfried E. Noether. Distribution-free confidence intervals. The American Statistician, 26(1):39–41, February 1972. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682846.

Noether:1977:BRB

[Noe77] Gottfried E. Noether. Book review: Didaktik der Mathematic. The American Statistician, 31(3):124, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682962.

Norris:1976:GMS

[Nor76] Nilan Norris. General means and statistical theory. The American Statistician, 30(1):8–12, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682879.

Norwood:1979:JS

[Nor79] Janet L. Norwood. Julius Shiskin, 1912–1978. The American Statistician, 33(3):137, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683816.

Novick:1975:CBS

[Nov75] Melvin R. Novick. A course in Bayesian statistics. The American Statistician, 29(2):94–98, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683274.

Norusis:1978:LE

[NVM^{+}78] Marija J. Norusis, Neal Van Eck, Richard G. Montanelli, Jr., Leland Wilkinson, Gerard E. Dallal, John Neter, John Petkau, and W. J. Conover. Letters to the Editor. The American Statistician, 32(3):113–114, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683362.

Oliver:1972:MLO

[Oli72] E. H. Oliver. A maximum likelihood oddity. The American Statistician, 26(3):43–44, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682867.

Orleans:1974:CSI

[Orl74] Leo A. Orleans. Chinese statistics: The impossible dream. The American Statistician, 28(2):47–52, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683598.

Padmanabhan:1977:ASW

[Pad77] A. R. Padmanabhan. Ancillary statistics which are not invariant. The American Statistician, 31(3):124, August 1977. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682961.

Panjer:1973:DMC

[Pan73] H. H. Panjer. On the decomposition of moments by conditional moments. The American Statistician, 27(4):170–171, October 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2684052.

Phillips:1974:LE

[PCM^{+}74] Haynes Phillips, P. Clifford, James H. Matis, Herbert F. Spirer, C. Terrence Ireland, John P. Hoyt, G. A. Barnard, Theodore D. Sterling, Samuel B. Lyerly, I. J. Good, Julian L. Simon, Dan Weidenfeld, V. P. Godambe, Ernest W. Grove, George A. F. Seber, Nathan Mantel, John E. Alman, T. W. Anderson, John A. Gaines, and Michael A. Golberg. Letters to the Editor. The American Statistician, 28(4):160–166, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683345.

Pirie:1979:LE

[PGG^{+}79] Walter R. Pirie, Michael L. Goodman, Hans U. Gerber, Jean D. Gibbons, Ingram Olkin, Milton Sobel, and David A. Pyne. Letters to the Editor. The American Statistician, 33(4):228–229, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683747.

Prescott:1977:TOM

[PH77] P. Prescott and R. V. Hogg. Trimmed and outer means and their variances. The American Statistician, 31(4):156–157, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683537.

Pieters:1976:SHS

[Pie76] Richard S. Pieters. Statistics in the high school curriculum. The American Statistician, 30(3):134–139, August 1976. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683860.

Paper:1979:LE

[PKT^{+}79] Elinor S. Paper, C. G. Khatri, Ajit C. Tamhane, David W. Smith, Nathan Mantel, and S. K. Lee. Letters to the Editor. The American Statistician, 33(2):92–94, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683233.

Pollock:1978:IRV

[Pol78] K. H. Pollock. Inference robustness vs. criterion robustness: An example. The American Statistician, 32(4):133–136, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2682940. See corri-

genda [Ano79f].

Posten:1974:ALB

[Pos74] Harry O. Posten. An abbreviated list of book reviews 1960–1973. The American Statistician, 28(4):138–142, November 1974. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683340.

Posten:1976:BAV

[Pos76] Harry O. Posten. A bibliography on audio-visual materials for statistical education. The American Statistician, 30(2):91–96, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683802.

Posten:1977:LH

[Pos77a] Harry O. Posten. Living histograms. The American Statistician, 31 (2):94, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683052.

Posten:1977:SBA

[Pos77b] Harry O. Posten. Supplement to a bibliography on audiovisual materials for statistical education. The American Statistician, 31 (4):163–165, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683540.

Posten:1978:RPS

[Pos78a] Harry O. Posten. Review of probability and statistics lab. The American Statistician, 32(3):108, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683358.

Posten:1978:RPD

[Pos78b] Harry O. Posten. Review of probability device. The American Statistician, 32(1):37, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683475.

Posten:1979:BRB

[Pos79] Harry O. Posten. Book review: [untitled]. The American Statistician, 33(3):148, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683824.

Prochaska:1973:NRB

[Pro73] B. J. Prochaska. A note on the relationship between the geometric and exponential distributions. The American Statistician, 27(1):27, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682902.

Pollock:1979:SGU

[PRPM79] K. H. Pollock, H. M. Ross-Parker, and R. Mead. A sequence of games useful in teaching experimental design to agriculture students. The American Statistician, 33(2):70–76, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683224.

Perlman:1975:SBN

[PW75] Michael D. Perlman and Michael J. Wichura. Sharpening Buffon’s needle. The American Statistician, 29(4):157–163, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683484.

Rubin:1971:QAc

[RA71] Ernest Rubin and Henry Anderson. Questions and answers. The American Statistician, 25(3):46–50, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683325.

Rubin:1973:QAa

[RA73] Ernest Rubin and Henry Anderson. Questions and answers. The American Statistician, 27(1):32–37, February 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682907.

Raff:1970:CGF

[Raf70] Morton S. Raff. On calculating the gamma function of non-integral arguments. The American Statistician, 24(2):22–24, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681926.

Rubin:1970:QAd

[RAM70] Ernest Rubin, William Alban, and Eleanor J. MacDonald. Questions and answers. The American Statistician, 24(4):31, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2681314.

Rao:1971:SNM

[Rao71] Potluri Rao. Some notes on misspecification in multiple regressions. The American Statistician, 25(5):37–39, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686082.

Rao:1973:SNE

[Rao73] Potluri Rao. Some notes on the errors-in-variables model. The American Statistician, 27(5):217–218, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/3087403.

Rao:1976:CBV

[Rao76] Potluri Rao. On a correspondence between t and F values in multiple regressions. The American Statistician, 30(4):190–191, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683761.

Rao:1977:NEW

[Rao77] C. Radhakrishna Rao. A natural example of a weighted binomial distribution. The American Statistician, 31(1):24–26, February 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682973.

Rubin:1970:QAa

[RB70] Ernest Rubin and Irwin D. J. Bross. Questions and answers. The American Statistician, 24(1):53–56, February 1970. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682301.

Rubin:1971:QAd

[RB71] Ernest Rubin and Henry S. Bashkin. Questions and answers. The American Statistician, 25(4):43–45, October 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682928.

Rabinowitz:1974:CVM

[RB74] M. Rabinowitz and M. L. Berenson. A comparison of various methods of obtaining random order statistics for Monte Carlo computations. The American Statistician, 28(1):27–29, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683528.

Rosenberg:1976:HRP

[RD76] W. J. Rosenberg and J. J. Deely. The horse-racing problem — a Bayesian approach. The American Statistician, 30(1):26–29, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682883.

Rubin:1971:QAa

[RGH71] Ernest Rubin, Robert Gillingham, and Dale Heien. Questions and answers. The American Statistician, 25(1):54–55, February 1971.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682219.

Rosenblatt:1970:PDA

[RGW70] David Rosenblatt, Ezra Glaser, and Marshall K. Wood. Principles of design and appraisal of statistical information systems. The American Statistician, 24(4):10–15, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681307.

Rice:1977:RSD

[Ric77] Dorothy P. Rice. The role of statistics in the development of health care policy. The American Statistician, 31(3):101–106, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682954.

Riddell:1977:PGL

[Rid77] Wm. Craig Riddell. Prediction in generalized least squares. The American Statistician, 31(2):88–90, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683049.

Ryan:1973:MSC

[RJ73] T. A. Ryan, Jr. and Brian L. Joiner. Minitab: a statistical computing system for students and researchers. The American Statistician, 27(5):222–225, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087405.

Roy:1974:PSE

[RLM74] Roch Roy, Yves Lepage, and Marc Moore. On the power series expansion of the moment generating function. The American Statistician, 28(2):58–59, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683601.

Rosenberg:1977:SDC

[RM77] Harry M. Rosenberg and George C. Myers. State demographic centers: Their current status. The American Statistician, 31 (4):141–146, November 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683534.

Roberts:1971:CSS

[Rob71] Harry V. Roberts. Committee selection by statistical sampling. The American Statistician, 25(1):18–20, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682206.

Roberts:1978:SCMa

[Rob78a] Harry V. Roberts. Statisticians can matter. The American Statistician, 32(2):45–51, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor. org/stable/2683612. See comments [Wel78, Win78] and rejoinder [Rob78b].

Roberts:1978:SCMb

[Rob78b] Harry V. Roberts. [statisticians can matter]: Rejoinder. The American Statistician, 32(2):57, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/ stable/2683615. See [Wel78, Win78, Rob78a].

Roberts:1979:IAT

[Rob79] Harry V. Roberts. [information and allocation: Two uses of the 1980 census]: Comment. The American Statistician, 33(2):52–54, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683220. See

[Key79a].

Rosander:1970:AKC

[Ros70] A. C. Rosander. Analysis of the Kaysen Committee Report. The American Statistician, 24(1):20–25, February 1970. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682291.

Rosenberg:1971:CDA

[Ros71] Harry M. Rosenberg. Census data and the analysis of social trends. The American Statistician, 25(5):24–27, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2686079.

Rosenblatt:1979:RSK

[Ros79] Judah Rosenblatt. Review of Statistical Kit No. 1. The American Statistician, 33(1):25, February 1979. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683063.

Rubin:1973:QAc

[RR73] Ernest Rubin and David Rothman. Questions and answers. The American Statistician, 27(4):186–191, October 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684054.

Rubin:1970:QAc

[RS70] Ernest Rubin and John H. Smith. Questions and answers. The American Statistician, 24(3):16–20, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682420.

Rubin:1972:QAc

[RS72] Ernest Rubin and John H. Smith. Questions and answers. The American Statistician, 26(3):59–61, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682870.

Rothman:1973:MPE

[RS73] David Rothman and William C. Shelton. Majorities and pluralities in elections. The American Statistician, 27(5):240, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/3087414.

Rubin:1970:QAb

[Rub70a] Ernest Rubin. Questions and answers. The American Statistician, 24(2):33–34, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681930.

Rubin:1970:QAe

[Rub70b] Ernest Rubin. Questions and answers. The American Statistician, 24(5):33–36, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682451.

Rubin:1971:QAb

[Rub71a] Ernest Rubin. Questions and answers. The American Statistician, 25(2):34–35, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682006.

Rubin:1971:QAe

[Rub71b] Ernest Rubin. Questions and answers. The American Statistician, 25(5):52–54, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2686088.

Rubin:1972:QAa

[Rub72a] Ernest Rubin. Questions and answers. The American Statistician, 26(1):42–44, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682847.

Rubin:1972:QAb

[Rub72b] Ernest Rubin. Questions and answers. The American Statistician, 26(2):34–36, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683461.

Rubin:1972:QAd

[Rub72c] Ernest Rubin. Questions and answers. The American Statistician, 26(4):44–47, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683212.

Rubin:1972:QAe

[Rub72d] Ernest Rubin. Questions and answers. The American Statistician, 26(5):37–39, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683786.

Rubin:1973:EMS

[Rub73a] Ernest Rubin. Estimating marriageable spouses in the single population. The American Statistician, 27(5):233–238, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/3087411.

Rubin:1973:QAb

[Rub73b] Ernest Rubin. Questions and answers. The American Statistician, 27(3):119–122, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683638.

Rust:1976:ELY

[Rus76] Philip F. Rust. The effect of leap years and seasonal trends on the birthday problem. The American Statistician, 30(4):197–198, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683764.

Rutherford:1971:NVC

[Rut71] J. R. Rutherford. A note on variances in the components of variance model. The American Statistician, 25(3):26–27, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683319.

Robertson:1974:CMD

[RW74] Tim Robertson and F. T. Wright. Consistency of the midrange: Distributions whose tails are both light and smooth. The American Statistician, 28(4):135–137, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683338.

Rich:1970:LE

[RWH^{+}70] Kenneth L. Rich, Edward J. Wegman, R. A. Hurtubise, David R. Segal, Raymond H. Burros, B. Eknath Rao, Joan M. Finucci, Julian C. Stanley, S. Mitra, S. R. Searle, A. L. Wilson, A. W. Douglas, Donald J. Wheeler, P. D. M. Macdonald, Charles E. Land, J. M. Maynard, and Alan J. Gross. Letters to the Editor. The American Statistician, 24(5):37–43, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682453.

Rathbun:1972:RPC

[RWM^{+}72] Daniel B. Rathbun, W. Allen Wallis, Frederick Mosteller, Ansley J. Coale, Paul M. Densen, Solomon Fabricant, Robert D. Fisher, W. Braddock Hickman, William Kruskal, Stanley Lebergott, Richard M. Scammon, William H. Shaw, Frank D. Stella, James A. Suffridge, and John W. Tukey. The report of the President’s Commission on Federal Statistics. The American Statistician, 26(1):15–18, February 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682840.

Ryan:1979:TSC

[Rya79] Barbara F. Ryan. [teaching statistical computing using computer packages]: Comment. The American Statistician, 33(1):33–34, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683068.

See [Thi79a].

Rybak:1975:DST

[Ryb75] Janet Rybak. Diagrams for set theory and probability problems of four or more variables. The American Statistician, 29(2):91–93, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683272.

Salsburg:1973:SWI

[Sal73] David S. Salsburg. Sufficiency and the waste of information. The American Statistician, 27(4):152–154, October 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684044.

Samaniego:1975:ME

[Sam75] Francisco J. Samaniego. On T -minimax estimation. The American Statistician, 29(4):168–169, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683487.

Sastry:1970:SLT

[Sas70] M. V. Rama Sastry. Some limits in the theory of multicollinearity. The American Statistician, 24(1):39–40, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682295.

Savage:1970:RSF

[Sav70] Leonard J. Savage. Reading suggestions for the foundations of statistics. The American Statistician, 24(4):23–27, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2681310.

Stoops:1971:MCC

[SB71] Glenn Stoops and Donald Barr. Moments of certain Cauchy order statistics. The American Statistician, 25(5):51, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2686087.

Selvin:1975:LE

[SBK^{+}75] Steve Selvin, M. Bloxham, A. I. Khuri, Michael Moore, Rodney Coleman, G. Rex Bryce, James A. Hagans, Thomas C. Chalmers, E. A. Maxwell, and Gary N. Smith. Letters to the Editor. The American Statistician, 29(1):67–71, February 1975. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683689.

Skvarcius:1971:NUC

[SC71] Romualdas Skvarcius and Fred Cromer. A note on the use of categorical vectors in testing for equality of two regression equations. The American Statistician, 25(3):27–29, June 1971. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683320.

Schwartz:1971:OTF

[Sch71] Ronald D. Schwartz. Operational techniques of a factor analysis model. The American Statistician, 25(4):38–42, October 1971. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682927.

Schafer:1976:MEB

[Sch76] Ray E. Schafer. Modified estimators for the binomial parameter. The American Statistician, 30(2):98–100, May 1976. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683804.

Scherer:1979:SGR

[Sch79a] F. M. Scherer. Statistics for governmental regulation. The American Statistician, 33(1):1–5, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683057.

Schulman:1979:GMR

[Sch79b] Robert S. Schulman. A geometric model of rank correlation. The American Statistician, 33(2):77–80, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683225.

Sclove:1972:MVE

[Scl72] Stanley L. Sclove. On missing value estimation in experimental design models. The American Statistician, 26(2):25–26, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683456.

Searle:1973:PEA

[Sea73] S. R. Searle. On publishing extended abstracts, and reviews. The American Statistician, 27(4):155–157, October 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2684045.

Searle:1979:ACO

[Sea79] S. R. Searle. Annotated computer output for analysis of variance of unequal-subclass-numbers data. The American Statistician, 33(4):222–223, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683742.

Seber:1973:INV

[Seb73] G. A. F. Seber. Independence and normal variables. The American Statistician, 27(3):110–111, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683633.

Seely:1977:ELH

[See77] Justus Seely. Estimability and linear hypotheses. The American Statistician, 31(3):121–123, August 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682960.

Seigel:1975:SAM

[Sei75] Daniel G. Seigel. Several approaches for measuring average rates of change for a second degree polynomial. The American Statistician, 29(1):36–37, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683678.

Selvin:1971:MLE

[Sel71] Steve Selvin. Maximum likelihood estimation in the truncated, single parameter, discrete exponential family. The American Statistician, 25(1):41–42, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682212.

Schawe:1977:LE

[SGM77] David Schawe, Gerald R. Garside, and Nathan Mantel. Letters to the Editor. The American Statistician, 31(3):134–135, August 1977.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682966.

Speed:1976:UNU

[SH76] F. M. Speed and R. R. Hocking. The use of the R()-notation with unbalanced data. The American Statistician, 30(1):30–33, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682884.

Shahabuddin:1979:TIC

[Sha79] Syed Shahabuddin. TIMESERIES — an interactive computer program for time series analysis. The American Statistician, 33(4):223, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683743.

Shelton:1972:MPE

[She72] William C. Shelton. Majorities and pluralities in elections. The American Statistician, 26(5):17–19, December 1972. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683774.

Sheehan:1978:SMS

[She78] T. Joseph Sheehan. Statistics for medical students: Personalizing the Keller plan. The American Statistician, 32(3):96–99, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683351.

Searle:1976:LE

[SHG76] S. R. Searle, J. J. Hubert, and Robert Gilchrist. Letters to the Editor. The American Statistician, 30(1):43–44, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682889.

Shiskin:1970:SFS

[Shi70] Julius Shiskin. Strengthening federal statistics. The American Statistician, 24(1):15–20, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682290.

Shiskin:1971:MBC

[Shi71a] Julius Shiskin. Modernizing business cycle concepts. The American Statistician, 25(4):17–19, October 1971. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682919.

Shiskin:1971:RDF

[Shi71b] Julius Shiskin. Recent developments in federal statistics. The American Statistician, 25(5):18–21, December 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2686077.

Shuster:1973:SMT

[Shu73] Jonathan Shuster. A simple method of teaching the independence of X̄ and s^{2}. The American Statistician, 27(1):29–30, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682904.

Shulte:1978:BRBb

[Shu78a] Albert P. Shulte. Book review: Proceedings of the Fourth ISI Round Table Conference. The American Statistician, 32(4):141–142, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682945.

Shulte:1978:BRBa

[Shu78b] Albert P. Shulte. Book review: Proceedings of the Third ISI Round Table Conference. The American Statistician, 32(4):140–141, November 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682944.

Siddiqui:1974:VLP

[Sid74] M. M. Siddiqui. The visiting lecturer program in statistics. The American Statistician, 28(3):94–96, August 1974. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683032.

Simons:1976:IAF

[Sim76] Gordon Simons. An interesting application of Fatou’s Lemma. The American Statistician, 30(3):146, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683863.

Sirotnik:1971:MMA

[Sir71] Ken Sirotnik. On the meaning of the mean in ANOVA (or the case of the missing degree of freedom). The American Statistician, 25(4):36–37, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682926.

Srirama:1971:LE

[SKR^{+}71] M. Srirama, Scott A. Krane, B. E. Rodden, Robert W. Kennard, Robert A. Agnew, Edward Pollak, I. J. Good, J. A. Hartigan, Edward E. Cureton, D. C. Crocker, C. Philip Cox, E. Ann Blair, David M. Brown, and Al Wilson. Letters to the Editor. The American Statistician, 25(3):52–56, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683328.

Sutherland:1979:BRB

[SL79] Mike Sutherland and Stanley Lemeshow. Book review: Numerian Population Simulator. The American Statistician, 33(3):145, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683821.

Speed:1979:ETM

[SM79] F. M. Speed and Charles J. Monlezun. Exact F tests for the method of unweighted means in a 2^{k} experiment. The American Statistician, 33(1):15–18, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683060.

Smith:1978:SAM

[Smi78] David W. Smith. A simplified approach to the maximum likelihood estimation of the covariance matrix. The American Statistician, 32(1):28–29, February 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683472.

Smith:1979:SUC

[Smi79] Patricia L. Smith. Splines as a useful and convenient statistical tool. The American Statistician, 33(2):57–62, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683222.

Snee:1974:GDT

[Sne74] Ronald D. Snee. Graphical display of two-way contingency tables. The American Statistician, 28(1):9–12, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683520.

Sogliero:1979:BRB

[Sog79] Gene Sogliero. Book review: [untitled]. The American Statistician, 33(3):147, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683823.

Solomon:1972:TS

[Sol72] Daniel L. Solomon. On testing for sufficiency. The American Statistician, 26(4):37, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683207.

Solomon:1975:NNE

[Sol75] Daniel L. Solomon. A note on the non-equivalence of the Neyman– Pearson and generalized likelihood ratio tests for testing a simple null versus a simple alternative hypothesis. The American Statistician, 29(2):101–102, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683277.

Speed:1974:AGI

[Spe74] F. M. Speed. An application of the generalized inverse to the oneway classification. The American Statistician, 28(1):16–18, February 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683522.

Sprinchorn:1970:OH

[Spr70] Evert Sprinchorn. The odds on Hamlet. The American Statistician, 24(5):14–17, December 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682444.

Searle:1974:NEC

[SR74] S. R. Searle and T. R. Rounsaville. A note on estimating covariance components. The American Statistician, 28(2):67–68, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683606.

Stewart:1970:NTL

[SS70] W. C. Stewart and B. R. Siskin. A non-typical least-squares solution to an engineering problem. The American Statistician, 24(1):47–48, February 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682298.

Sampson:1976:SMS

[SS76] Allan Sampson and Bruce Spencer. Sufficiency, minimal sufficiency, and the lack thereof. The American Statistician, 30(1):34–35, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682885.

See corrigenda [Ano77e].

Sahai:1973:NNM

[ST73] Hardeo Sahai and William O. Thompson. Non-negative maximum likelihood estimators of variance components in a simple linear model. The American Statistician, 27(3):112–113, June 1973. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683634.

Steele:1973:WSF

[Ste73a] J. M. Steele. When successes and failures are independent a compound process is Poisson. The American Statistician, 27(5):232, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/3087410.

Sterling:1973:SVV

[Ste73b] Theodor D. Sterling. The statistician vis-a-vis issues of public health. The American Statistician, 27(5):212–217, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/3087402.

Stigler:1972:CUE

[Sti72] Stephen M. Stigler. Completeness and unbiased estimation. The American Statistician, 26(2):28–29, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683458.

Strait:1971:NIC

[Str71] Peggy Tang Strait. A note on independence and conditional probabilities. The American Statistician, 25(2):17–18, April 1971. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682000.

Strand:1979:EPT

[Str79] Michael M. Strand. Estimation of a population total under a “Bernoulli Sampling” procedure. The American Statistician, 33(2): 81–84, May 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683226.

Sweeney:1972:TSI

[SU72] Robert E. Sweeney and Edwin F. Ulveling. A transformation for simplifying the interpretation of coefficients of binary variables in regression analysis. The American Statistician, 26(5):30–32, December 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683780.

Swindel:1974:IRC

[Swi74] Benee F. Swindel. Instability of regression coefficients illustrated. The American Statistician, 28(2):63–65, May 1974. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683604.

Simonoff:1979:LE

[SZM79] Jeffrey Simonoff, Alexis Zinger, and Gary M. Mullet. Letters to the Editor. The American Statistician, 33(1):36–37, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683071.

Tanis:1979:ICB

[Tan79] Elliot A. Tanis. An instructional computer-based package for probability and statistics. The American Statistician, 33(4):224–225, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683745.

Taubman:1972:BTH

[Tau72] Stephen Taubman. Best k of 2k − 1 tests of hypotheses involving the normal distribution. The American Statistician, 26(4):32–33, October 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683205.

Tiao:1973:SCB

[TB73] G. C. Tiao and G. E. P. Box. Some comments on “Bayes” estimators. The American Statistician, 27(1):12–14, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682897.

Temkin:1973:SSP

[Tem73] Sanford Temkin. Statistics and social problems. The American Statistician, 27(5):245–246, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/3087423.

Tenenbein:1971:RCP

[Ten71] Aaron Tenenbein. The racing car problem. The American Statistician, 25(1):38–40, February 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682211.

Thisted:1979:TSCa

[Thi79a] Ronald A. Thisted. Teaching statistical computing using computer packages. The American Statistician, 33(1):27–30, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683065. See

comments [DuM79, Hei79, Rya79, Wel79] and rejoinder [Thi79b].

Thisted:1979:TSCb

[Thi79b] Ronald A. Thisted. [teaching statistical computing using computer packages]: Rejoinder. The American Statistician, 33(1):35, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683070. See

[Thi79a].

Tarter:1976:IIT

[TK76] M. E. Tarter and R. A. Kronmal. An introduction to the implementation and theory of nonparametric density estimation. The American Statistician, 30(3):105–112, August 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683854.

Tortora:1978:NSS

[Tor78] Robert D. Tortora. A note on sample size estimation for multinomial populations. The American Statistician, 32(3):100–102, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683352.

Trawinski:1978:BRB

[Tra78] Irene M. Trawinski. Book review: Distribution Series. The American Statistician, 32(4):142–143, November 1978. CODEN ASTAAJ.

ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682946.

Urquhart:1971:NIA

[Urq71] N. S. Urquhart. Nonverbal instructional approaches for introductory statistics. The American Statistician, 25(2):20–25, April 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682002.

Urquhart:1975:SSC

[Urq75] N. Scott Urquhart. Segmenting service courses. The American Statistician, 29(1):47–50, February 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683683.

Varian:1972:LEB

[Var72] Hal R. Varian. Letter to the Editor: Benford’s Law. The American Statistician, 26(3):65–66, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:/

/links.jstor.org/sici?sici=0003-1305%28197206%2926%3A3%
3C62%3ALTTE%3E2.0.CO%3B2-Q.

VanMatre:1976:SEW

[VC76] Joseph G. Van Matre and William N. Clark. The statistician as expert witness. The American Statistician, 30(1):2–5, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682878.

VanMatre:1979:RTM

[VK79] Joseph G. Van Matre and Alan H. Kvanli. Review of teaching materials on census data. The American Statistician, 33(4):211–212, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2683737.

Wackerly:1976:DCS

[Wac76] Dennis D. Wackerly. On deriving a complete sufficient statistic. The American Statistician, 30(1):37–38, February 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682887.

Wainer:1974:SRO

[Wai74] Howard Wainer. The suspended rootogram and other visual displays: An empirical validation. The American Statistician, 28 (4):143–145, November 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683341.

Wallenius:1971:CCF

[Wal71] K. T. Wallenius. A conditional covariance formula with applications. The American Statistician, 25(3):32–33, June 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683323.

Wallis:1976:SNC

[Wal76] W. Allen Wallis. Statistics in nonstatistical contexts. The American Statistician, 30(4):159–164, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683752.

Warren:1973:PC

[War73] W. G. Warren. On partial correlation. The American Statistician, 27(5):239, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/3087412.

Watts:1973:TMT

[Wat73] Donald G. Watts. Transformations made transparently easy, or, so that’s what a Jacobian is! The American Statistician, 27(1):22–25, February 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-

2731 (electronic). URL http://www.jstor.org/stable/2682900.

Walster:1970:PNE

[WC70] G. William Walster and T. Anne Cleary. A proposal for a new editorial policy in the social sciences. The American Statistician, 24(2):16–19, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2681924.

Wilkinson:1977:ASM

[WD77] Leland Wilkinson and Gerard E. Dallal. Accuracy of sample moments calculations among widely used statistical programs. The American Statistician, 31(3):128–131, August 1977. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2682964.

Weiss:1970:GRR

[Wei70] Neil S. Weiss. A graphical representation of the relationships between multiple regression and multiple correlation. The American Statistician, 24(2):25–29, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2681927.

Weiner:1974:RPC

[Wei74] Neil S. Weiner. The report of the President’s Commission on Federal Statistics: a summary of recommendations, views and counterviews. The American Statistician, 28(2):42–46, May 1974. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683597.

Wells:1975:MJB

[Wel75] Oris V. Wells. Mordecai J. B. Ezekiel, 1899–1974. The American Statistician, 29(2):106, May 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683279.

Welsch:1978:SCM

[Wel78] Roy E. Welsch. [statisticians can matter]: Comment. The American Statistician, 32(2):52–53, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/ stable/2683613. See [Rob78a, Win78, Rob78b].

Welsch:1979:TSC

[Wel79] Roy E. Welsch. [teaching statistical computing using computer packages]: Comment. The American Statistician, 33(1):34, February 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2683069. See

[Thi79a].

Wertheimer:1977:LE

[WER^{+}77] Isabel Wertheimer, A. S. C. Ehrenberg, Herbert Robbins, David C. Hoaglin, R. W. Farebrother, and John W. Pratt. Letters to the Editor. The American Statistician, 31(2):97–99, May 1977. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683054.

Wolfe:1971:CSR

[WH71] Douglas A. Wolfe and Robert V. Hogg. On constructing statistics and reporting data. The American Statistician, 25(4):27–30, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2682922.

White:1971:CP

[Whi71] Colin White. The committee problem. The American Statistician, 25(4):25–26, October 1971. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682921.

Whittaker:1973:VDS

[Whi73] J. Whittaker. A visual derivation of some elementary discrete distributions. The American Statistician, 27(3):116–118, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683637.

Wickens:1979:IL

[Wic79] Aryness Joy Wickens. Isador Lubin, 1896–1978. The American Statistician, 33(4):207, November 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683734.

Willis:1973:CPC

[Wil73a] Raymond E. Willis. Confidence procedures and the cost of sampling. The American Statistician, 27(5):219–221, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/3087404.

Wilson:1973:ZCI

[Wil73b] William J. Wilson. Zero covariance and independence. The American Statistician, 27(5):242, December 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/3087417.

Willits:1976:JSD

[Wil76] Joseph H. Willits. Joseph Stancliffe Davis, (1885–1975). The American Statistician, 30(4):199, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683765.

Williams:1978:HBC

[Wil78a] W. H. Williams. How bad can “good” data really be? The American Statistician, 32(2):61–65, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683617.

Wilson:1978:WCT

[Wil78b] A. L. Wilson. When is the Chow test UMP? The American Statistician, 32(2):66–68, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/ stable/2683618. See corrigenda [Ano79h].

Winkler:1978:SCM

[Win78] Robert L. Winkler. [statisticians can matter]: Comment. The American Statistician, 32(2):54–56, May 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www. jstor.org/stable/2683614. See [Rob78a, Wel78, Rob78b].

Winterbottom:1979:NDF

[Win79] Alan Winterbottom. A note on the derivation of Fisher’s transformation of the correlation coefficient. The American Statistician, 33(3):142–143, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683819.

Wiorkowski:1972:CAK

[Wio72] John J. Wiorkowski. A curious aspect of knockout tournaments of size 2^{m}. The American Statistician, 26(3):28–30, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682862.

Wiorkowski:1975:WCI

[Wio75] John J. Wiorkowski. The Wind Chill Index — a case study in table sleuthing. The American Statistician, 29(4):152–156, November 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683483.

Wani:1970:CMP

[WK70] J. K. Wani and D. G. Kabe. On a certain minimization problem. The American Statistician, 24(4):29–30, October 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681313.

Wagner:1972:PSI

[WM72] G. R. Wagner and B. Motazed. The proctorial system of instruction combined with computer pedagogy for teaching applied statistics. The American Statistician, 26(3):36–39, June 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2682865.

Wagner:1973:ICQ

[WM73] G. R. Wagner and M. M. McCants. An interactive computer quizzing system to support the proctorial system of instruction. The American Statistician, 27(3):113–115, June 1973. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2683635.

Wong:1972:NMI

[Won72] C. K. Wong. A note on mutually independent events. The American Statistician, 26(2):27–28, April 1972. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683457.

Woodworth:1976:TPA

[Woo76] George G. Woodworth. t for two, or preposterior analysis for two decision makers: Interval estimates for the mean. The American Statistician, 30(4):168–171, November 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683754.

Weiss:1970:LE

[WRH^{+}70] Moshe Weiss, J. N. K. Rao, H. O. Hartley, W. Edwards Deming, David Durand, Edward J. Dudewicz, Mrs. Maryeve Schick, Hans K. Ury, James C. Forrester, Andreas N. Philippou, Ram C. Dahiya, Potluri Rao, Henry E. Schaffer, Thomas F. Green, Prakash C. Joshi, and Morton S. Raff. Letters to the Editor. The American Statistician, 24(3):20–29, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2682421.

Wolkowicz:1979:ESI

[WS79] Henry Wolkowicz and George P. H. Styan. Extensions of Samuelson’s inequality. The American Statistician, 33(3):143–144, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.jstor.org/stable/2683820.

Wadycki:1973:LE

[WSG^{+}73] Walter J. Wadycki, B. K. Shah, P. D. Ghangurde, Edward J. Dudewicz, Nathan Mantel, Charles C. Brown, Harold J. Larson, Donald R. Barr, James W. Frane, Bernard Saperstein, I. J. Good, and Howard L. Jones. Letters to the Editor. The American Statistician, 27(3):123–127, June 1973. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2683639.

Yeh:1970:PGI

[Yeh70] Chiao Yeh. Probability and the game of instant insanity. The American Statistician, 24(3):13, June 1970. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2682418.

Young:1979:WRB

[You79] Ralph A. Young. Warren Randolph Burgess, 1889–1978. The American Statistician, 33(3):136, August 1979. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683815.

Young:1975:ERI

[YY75] Kan Hua Young and Lin Ying Young. Estimation of regressions involving logarithmic transformation of zero values in the dependent variable. The American Statistician, 29(3):118–120, August 1975. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2683436.

Young:1976:LE

[YYB^{+}76] Kan Hua Young, Lih Ying Young, A. E. Brandt, C. E. Werts, D. A. Rock, R. L. Linn, K. C. Joreskog, R. W. Farebrother, Oliver D. Anderson, Federico J. O’Reilly, J. A. Nelder, Nathan Mantel, Mark J. Nicolich, and John Neter. Letters to the Editor. The American Statistician, 30(2):101–104, May 1976. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2683805.

Zerbe:1978:FTG

[Zer78] Gary O. Zerbe. On Fieller’s theorem and the general linear model. The American Statistician, 32(3):103–105, August 1978. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2683354.